### v2

# Statistical Inferences and Predictions for Areal Data and Spatial Data Fusion with Hausdorff–Gaussian Processes

Abstract
Accurate modeling of spatial dependence is pivotal in analyzing spatial data, influencing parameter estimation and out-of-sample predictions. The spatial structure and geometry of the data significantly impact valid statistical inference. Existing models for areal data often rely on adjacency matrices, struggling to differentiate between polygons of varying sizes and shapes. Conversely, data fusion models, while effective, rely on computationally intensive numerical integrals, presenting challenges for moderately large datasets. In response to these issues, we propose the Hausdorff-Gaussian process (HGP), a versatile model class utilizing the Hausdorff distance to capture spatial dependence in both point and areal data. We introduce a valid correlation function within the HGP framework, accommodating diverse modeling techniques, including geostatistical and areal models. Integration into generalized linear mixed-effects models enhances its applicability, particularly in addressing change of support and data fusion challenges. We validate our approach through a comprehensive simulation study and application to two real-world scenarios: one involving areal data and another demonstrating its effectiveness in data fusion. The results suggest that the HGP is competitive with specialized models regarding goodness-of-fit and prediction performances. In summary, the HGP offers a flexible and robust solution for modeling spatial data of various types and shapes, with potential applications spanning fields such as public health and climate science.
Keywords: Areal data, CAR models, Geostatistics, Spatial modeling, Spatial prediction
1 Introduction
Accounting for spatial dependence is pivotal in spatial data analysis. Valid statistical inference are deeply influenced by the spatial structure and geometry of the data. Point-referenced data are typically analyzed using geostatistics methods (e.g., Cressie, 1993). For areal data, the conditional autoregressive model (CAR) (Besag, 1974) and its variants (Besag et al., 1991; Rodrigues and Assunçao, 2012; Datta et al., 2019; Cruz-Reyes et al., 2023) are often preferred. The nature of the observed spatial data significantly influences the choice of statistical methodology. This becomes increasingly critical in complex inferential problems, such as predicting a variable in a target map using data from a source map. This issue, known as change of support (Cressie, 1993, pg. 284–289), requires tailored methods depending on the spatial units involved (Gelfand et al., 2001). Spatial data from diverse sources often differ in spatial units or supports, resulting in spatial misalignment (Gotway and Young, 2002). Data fusion is necessitated when a spatial variable is provided by multiple, potentially misaligned, sources (Moraga et al., 2017).
Both traditional areal models and data fusion models for spatial data exhibit inherent limitations. Traditional areal models face challenges with predictions for overlapping spatial units (Bakar and Jin, 2020), do not account for region size or shape of the areal units and lack interpretable spatial dependence parameters (Assunção and Krainski, 2009; Cruz-Reyes et al., 2023). An exception of the last limitation is the directed acyclic graph auto-regressive (DAGAR) model, which offers an interpretable parameter (Datta et al., 2019). Data fusion models, which view areal data as averages of a continuous process, may be unsuitable for discrete outcomes (Gelfand et al., 2001). Additionally, their reliance on arbitrary discretization for likelihood evaluation lacks consensus on optimal resolution (Fuentes and Raftery, 2005; Liu et al., 2011; Wang and Furrer, 2021) and can introduce unquantifiable biases (Gonçalves and Gamerman, 2018).
We propose a unified approach for analyzing spatial data encompassing areal, point-referenced, and mixed spatial supports. This methodology combines Gaussian processes (GPs) with the Hausdorff distance (Shephard and Webster, 1965). The Hausdorff distance, defined on a reference metric space, effectively measures between the distance between two sets while accounting for their shape and size. For point pairs, it reduces to the distance metric associated to the reference metric space. GPs, recognized for their flexibility, are widely utilized in spatial data modeling, notably as priors in generalized mixed effects models (Diggle et al., 1998), for spatially varying regression coefficients (Gelfand et al., 2003), and in addressing spatial misalignment issues (Banerjee and Gelfand, 2002). The Hausdorff–Gaussian process (HGP) that we propose is a GP with a correlation function defined on the Hausdorff distance between sample units, allowing integrating their shape and size into the analysis. With some ensurance of positiveness definiteness of the resulting covariance matrix, the HGP with a powered exponential correlation can be used in place of GPs to adeptly model areal, point-referenced, and mixed spatial data types, offering a new avenue in analyzing spatial and spatiotemporal data with mixed supports.
This paper is structured as follows. Section 2 introduces the Hausdorff distance, discussing its properties, and then delineates the HGP, highlighting its significant applications. Section 3 sets up spatial generalized linear mixed models (GLMM) with HGP components and details their inferences in the Bayesian framework. A simulation study to compare the proposed model with specific models for areal and fused data is presented in Section 4. Section 5 demonstrates the practical utility of our model through two distinct applications, one on areal data and the other on data fusion. Section 6 concludes with a discussion.
2 Hausdorff–Gaussian Process
2.1 Hausdorff Distance
The Hausdorff distance is a metric measuring the distance between two sets of points. Specifically, it is the greatest distance from a point in one set to the closest point in the other set. Consider a metric space $(D,d)$, where $D$ is a spatial region of interest and $d(x,y)$ defines the distance between any two elements $x,y\in D$. For a single point $x$ and a set $A$, define $d(x,A)=\inf_{a\in A}d(x,a)$ as the closest distance from $x$ to any element in $A$. The directed Hausdorff distance from a set $A$ to a set $B$ is $\vec{h}(A,B)=\sup_{a\in A}d(a,B).$ This measure is not necessarily symmetric. The symmetric Hausdorff distance is defined as the greater of the two directed Hausdorff distances: $h(A,B)=\max\left\{\vec{h}(A,B),\vec{h}(B,A)\right\}.$ Note that if $A$ and $B$ are both singletons, then $h(A,B)=d(A,B)$.
In spatial analysis, the Hausdorff distance provides a robust and suitable metric over non-empty, bounded, and compact subsets of $D$, denoted as $\mathcal{B}(D)$ (Sendov, 2004). While not functioning as a conventional distance in unbounded sets, it is highly effective in bounded contexts, which are prevalent in spatial statistics. Typically, the study region $D$ in spatial statistics is a bounded subset of $\mathbb{R}^{2}$, where the Hausdorff distance excels in quantifying distances between spatial units by capturing both their shapes and sizes.
AB
AB
A.B
Figure 1: Three different pairs of bounded subsets of the Euclidean plane. In (a) the sets AA and $B$ share a single point on their border. In (b), they have non-empty intersection. While in (c), $B$ is a singleton set contained by $A$. The dashed lines represent the Hausdorff distance between the sets.
To demonstrate the efficacy of the Hausdorff distance in measuring distances between various spatial units, Figure 1 presents three simple examples. In each sub-figure, set $A$ is represented as a circle with a radius of 1. Figure 1(a) introduces a set $B$, a circle tangent to $A$. Figure 1(b) examines the distance between two sets having a non-empty intersection. Figure 1(c) features $B$ as an interior point of $A$. The dashed lines in the figures delineate the Hausdorff distances between the sets, highlighting the metric’s ability to consider the shape, size, and orientation of the objects. This contrasts with other simpler distance functions between non-empty sets, such as the infimum distance between two sets, denoted $\inf_{x\in X}\inf_{y\in Y}d(x,y)$. This distance does not fulfill the metric property of positiveness condition. In other words, $d(x,y)=0$ does not imply $x=y$. Indeed, employing such distance latter would yield a distance of zero between the three pairs of sets depicted in Figure 1.
2.2 Hausdorff–Gaussian Process
HGP is a natural extension of GP for modeling spatial data defined on the Hausdorff distance. A GP is a stochastic process over an index set such that all its finite-dimensional marginal distributions are multivariate normal distributions. An HGP expands the index set on a spatial region $D$ to encompasses a broader range of spatial units, $\mathcal{B}(D)$, which contains non-empty, bounded, and compact subsets of $D$. To define an isotropic HGP $Z(\cdot)$ over $\mathcal{B}(D)$, we specify a marginal mean function function $m(\cdot)$, a positive-valued marginal standard deviation (SD) function $v(\cdot)$, and an isotropic correlation function $r(h)$ for any two sets apart in Hausdorff distance $h$. With these functions, an HGP over $\mathcal{B}(D)$ is a stochastic process $Z(\cdot)$ whose marginal distribution at $\mathbf{s}_{1},\ldots,\mathbf{s}_{p}\in\mathcal{B}(D)$ is multivariate normal with mean vector and covariance matrix, respectively,
$\mathbbm{E}[Z(\mathbf{s}_{i})]=m(\mathbf{s}_{i})\quad\text{and}\quad\mathrm{Cov}[Z(\mathbf{s}_{i}),Z(\mathbf{s}_{j})]=v(\mathbf{s}_{i})v(\mathbf{s}_{j})r\big(h(\mathbf{s}_{i},\mathbf{s}_{j})\big),$ (1)
for $i,j=1,\ldots,p$, where $h(\mathbf{s}_{i},\mathbf{s}_{j})$ is the Hausdorff distance between $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$. This process allows for representations of diverse spatial data, from areal to mixed, making it suitable for a wide range of applications such as disease mapping and data fusion.
Denote the HGP process defined in Equation (1) by $\textrm{HGP}\{m(\cdot),v(\cdot),r(\cdot)\}$. The mean function $m(\cdot)$ captures large-scale variations, usually through covariates. The SD function $v(\cdot)$ function accommodates both homoscedastic and hereroscedastic scenarios. For instance, with a log link and covariate $w(\mathbf{s})$ at spatial unit $\mathbf{s}$, a natural model is
$\log v(\mathbf{s})=\alpha_{0}+\alpha_{1}w(\mathbf{s}).$ (2)
Equation (2) allows a distinction between point-referenced and areal data, a need in data fusion applications (Wang and Furrer, 2021). With $\mathcal{A}(\cdot)$ being the function that returns the area of a spatial object. A covariate $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$ enables different variances between these data types. Certainly, care is necessary in specifying $v(\mathbf{s})$ to ensure a valid GP (Palacios and Steel, 2006).
2.3 Correlation Function
The correlation function $r(\cdot)$ needs to be carefully chosen to ensure the validity of the HGP as it defines the spatial dependence structure across different locations. In particular, we focus on isotropic correlation functions. Therefore, unless explicitly mentioned, all subsequent references to correlation functions will assume they are isotropic. A valid correlation function must be positive definite and meet two key criteria: $r(0)=1$ and $\lim_{h\to\infty}r(h)=0$ (Cressie, 1993, Ch. 2.3.2), where $h$ is the Hausdorff distance between a pair $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{B}(D)$. The latter ensures that spatial correlation diminishes with increasing distance. It is noteworthy that while most geostatistical correlation functions are constructed for Euclidean spaces (Cressie, 1993, Ch. 2.5.1), their positive definiteness may not extend to non-Euclidean spaces (Curriero, 2006; Li et al., 2023). This raises the question of whether existing frameworks can be adapted or if entirely new formulations are needed to maintain the mathematical rigor in non-Euclidean settings.
For practical purposes, we propose to use a two-parameter powered exponential correlation (PEC) function
$r(h;\varphi,\nu)=\exp\left\{-\frac{h^{\nu}}{\varphi^{\nu}}\right\},$
where $\nu$ controls the smoothness and $\varphi$ governs the range of the spatial dependence. To enhance interpretability, we reparametrize this function with $\rho={\log(10)}^{1/\nu}\varphi$. Consequently, $\rho$ is the distance at which the spatial correlation between a pair of spatial units drops to $0.10$, commonly referred to as the practical range. The PEC function is appealing because the following Proposition 2.1 guarantees its positive definiteness for $\nu\in(1/2,1)$. The proof of the proposition is based on probabilistic embedding of Hausdorff metrics (Backurs and Sidiropoulos, 2016).
Proposition 2.1.
Let $h=h(\mathbf{s},\mathbf{s}^{\prime})$ be the Hausdorff distance between two spatial units, denoted $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{B}(D)$, where $D\subset\rm I\!R^{2}$. The powered exponential correlation function $\exp\{-h^{\nu}/\varphi^{\nu}\}$ is positive definite for $\nu\in(1/2,1)$.
In addition to the theoretical result, an empirical assessment was used to provide evidence for the positive definiteness of this function for an expanded range of $\nu\in(0,1)$. The positive definiteness is satisfied for bigger $\nu$ values when the dependence range $\rho$ is smaller. See details in Section LABEL:sec:sm:emp of the appendix.
3 HGP in Spatial GLMM
3.1 Model
Generalized linear models for spatial data can benefit from replacing the GP with the HGP as spatial random effects. In this context, assume $\mathcal{S}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}$ is a set of spatial units or locations such that $\mathbf{s}_{i}\in\mathcal{B}(D)$, where $D\subset\rm I\!R^{2}$. Consider a response variable $Y(\mathbf{s}_{i})$ and a $p$-dimensional vector of covariates $\mathbf{X}_{i}$ associated to the $i$-th spatial unit. Let $\mathbf{Z}={\big[Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n})\big]}^{\top}$ be an $n$-dimensional vector of latent spatial random effects. Conditional on $Z(\mathbf{s}_{i})$ and $\mathbf{X}_{i}$, we assume that the responses $Y(\mathbf{s}_{i})$’s are mutually independent. The conditinal distribution of $Y(\mathbf{s}_{i})$ is characterized by its mean mean and possibly some additional parameters. This leads to a spatial GLMM built on an HGP.
Specifically, the spatial GLMM is specified as
$Y(\mathbf{s}_{i})\mid\mathbf{X}_{i},Z(\mathbf{s}_{i})\overset{{\rm ind}}{\sim}f(\cdot\mid\mu_{i},\bm{\gamma}),$ (3)
where $f(\cdot\mid\mu_{i},\bm{\gamma})$ is a probability density or mass function of a distribution with expectation $\mu_{i}$ and an additional parameter vector $\bm{\gamma}$. The conditional mean $\mu_{i}$ is linked to the covariate vector and the spatial random effects through a suitable, known link function $g$:
$\mu_{i}=\mathbbm{E}[Y(\mathbf{s}_{i})\mid\mathbf{X}_{i},Z(\mathbf{s}_{i})]=g^{-1}(\mathbf{X}^{\top}_{i}\bm{\beta}+Z(\mathbf{s}_{i})),$ (4)
where $\bm{\beta}$ is a $p$-dimensional regression coefficient vector. The spatial dependence among the responses $\mathbf{Y}={\big[Y(\mathbf{s}_{1}),\ldots,Y(\mathbf{s}_{n})\big]}^{\top}$ is introduced by an HGP on $\mathbf{Z}$, namely,
$\mathbf{Z}\sim\textrm{HGP}\big\{0,v(\cdot;\bm{\sigma}),r(\cdot;\bm{\delta})\big\},$ (5)
with zero-mean, a SD function $v(\cdot;\bm{\sigma})$ with parameters $\bm{\sigma}$, and the PEC function $r(\cdot,\bm{\delta})$ with parameters $\bm{\delta}=\{\rho,\nu\}$. To summarize, the parameters of the spatial GLMM are $\bm{\theta}={\{\bm{\beta}^{\top},\bm{\sigma}^{\top},\bm{\delta}^{\top},\bm{\gamma}^{\top}\}}^{\top}$.
A general spatial GLMM is obtained by choosing a suitable $f(\cdot\mid\mu_{i},\bm{\gamma})$ for the response. For binary response variables, a Bernoulli distribution yields a spatial logistic regression. Count data can be modeled with either a Poisson or Negative Binomial distribution, depending on the presence of overdispersion. When the conditional distribution of the response is normal, the additional parameter $\gamma$ is the measurement error or small scale variance parameter. Let $\mathbf{X}$ be a $n\times p$ matrix with its $i$-th row given by $\mathbf{X}_{i}$. From the conditional dependence, the joint density (or mass) function of $\mathbf{Y}$ conditioning on $\mathbf{X}$ and $\mathbf{Z}$ is
$p(\mathbf{y}\mid\mathbf{X},\mathbf{z})=\prod_{i=1}^{n}f(y(\mathbf{s}_{i})\mid\mu_{i},\bm{\gamma}),$ (6)
where $\mathbf{y}$ and $\mathbf{z}$ are realizations of $\mathbf{Y}$ and $\mathbf{Z}$, respectively.
3.2 Bayesian Inferences
Prior distributions for all the model parameters are to be specified to complete the Bayesian model specification. For fixed effects regression coefficients ($\bm{\beta}$), we employ uncorrelated zero-mean Normal priors with marginal variances $10^{2}$. In the homoscedastic case, on $\sigma$ we place a half-t prior with $3$ degrees of freedom, denoted $t_{+}(3)$ (Gelman, 2006). For the hereroscedastic model (2), we set independent standard normal priors for $\alpha_{0}$ and $\alpha_{1}$. The PEC dependence is encoded through $\bm{\delta}=\{\rho,\nu\}$. In geostatistics, smoothness parameters for correlation functions are typically fixed, and their estimation properties are rarely reported (Zhang, 2004). Due to this and the scope of this paper, we specify separate models for different values of $\nu$ ensuring a positive definite PEC and select the best fit for inference. The spatial dependence parameter $\rho$ is also inherently difficult. Thus, to avoid oversmoothing, we employ an exponential prior for $\rho$ with mean $\lambda_{\rho}=-\rho_{0}/\log(p_{\rho})$, where $\rho_{0}$ is an upper-bound set to four-fifths of the largest observed Hausdorff distance, and $p_{\rho}$ is the probability that $\rho$ exceeds $\rho_{0}$. The resulting prior concentrates its mass on lower values $\rho$.
The posterior distribution of $\bm{\theta}$ and $\mathbf{Z}$ given the observed data $\mathbf{y}$ and $\mathbf{X}$ is
$\pi(\bm{\theta},\mathbf{z}\mid\mathbf{y})\propto p(\mathbf{y}\mid\mathbf{X},\mathbf{z},\bm{\beta},\bm{\gamma})\pi(\mathbf{z}\mid\bm{\sigma},\bm{\delta})\pi(\bm{\theta}),$ (7)
where $\pi(\cdot)$ denotes prior and posterior distributions of its arguments. Importantly, the distribution placed on the random effects $\mathbf{z}$ also functions as a prior, as we draw samples from these latent effects distribution conditioned on the observed data.
Parameter Estimation
The estimation of the spatial GLMM parameters is obtained from the posterior distribution defined in Equation (7). As the posterior distribution is intractable, we use the No-U-Turn (Hoffman and Gelman, 2014) Markov Chain Monte Carlo (MCMC) sampler available in Stan (Team, 2023) to draw samples from it. The No-U-Turn algorithm samples all the model parameters jointly and efficiently exploits the parameter space using automatic differentiation. Additionally, it eliminates the need for hand-tuning, making it a highly convenient sampler. We initialize the parameters with samples from their respective prior distributions. The latent random effects are initialized from a standard Gaussian distribution. The number of samples and the warm-up period of the MCMC algorithm are application dependent. We assess the convergence of the chains using the split-$\hat{R}$ diagnostic (Vehtari et al., 2021). Finally, the parameters’ point estimates and 95% credible intervals (CI) are obtained as, respectively, the median and percentiles ($2.5$ and $97.5$, unless otherwise stated) of the marginal MCMC samples.
Model Comparison
The assessment of goodness-of-fit (GoF) in the simulation study and applications involved two model comparison criteria: the leave-one-out information criterion (Vehtari et al., 2017, LOOIC), and the widely applicable information criterion (Watanabe, 2010, WAIC). For both of them, lower values indicate a better fit. We did not use the deviance information criterion due to its assumption of a small number of parameters relative to the sample size (McElreath, 2016, Ch. 7.4), which would require computationally intensive numerical integration of random effects in some of our models. When comparing a single candidate model to the HGP, we utilize the difference in the GoF metric $G$ between the HGP and the given model, denoted by $\Delta_{G}$. Consequently, a negative $\Delta_{G}$ suggests that the HGP model is a better fit for the data. The competing methods depend on the type of application we are working with and will be mentioned in the upcoming sections.
Predictions
Predicting spatial phenomena accurately at unobserved locations is crucial in many applications, especially in change of support settings (Gelfand et al., 2001). The HGP prior for spatial random effects in GLMMs offers a powerful solution. For a set of $m$ unobserved locations $\mathcal{S}^{\ast}=\{\mathbf{s}^{\ast}_{1},\ldots,\mathbf{s}^{\ast}_{m}\}$, we define an $m$-dimensional variable $\mathbf{Y}^{\ast}$ representing the response variable at these locations. Assuming covariates from Equation (4) are available, we can generate predictions accounting for uncertainty using the posterior predictive distribution of $\mathbf{Y}^{\ast}$:
$p(\mathbf{y}^{\ast}\mid\mathbf{y})=\int p(\mathbf{y}^{\ast}\mid\mathbf{z}^{\ast},\bm{\theta})p(\mathbf{z}^{\ast}\mid\mathbf{z},\bm{\theta})\pi(\bm{\theta}\mid\mathbf{y}){\rm d}\bm{\theta},$ (8)
where $\mathbf{z}^{\ast}={(z(\mathbf{s}^{\ast}_{1}),\ldots,z(\mathbf{s}^{\ast}_{m}))}^{\top}$ is a realization of the spatial random effect at the new locations, and $\bm{\theta}$ represents the model parameters. Details on sampling from this distribution are provided in Section LABEL:sec:sm:hgp-pred of the appendix.
4 Simulation Studies
4.1 Areal Data
We conducted a simulation study to assess the performance of the HGP relative to the DAGAR model, a state-of-the-art model designed explicitly for areal data (Datta et al., 2019), in modeling data generated from other models. We grounded our simulations in a real-world context by utilizing the 134 Intermediate Zones (IZ) north of the River Clyde within the Greater Glasgow in Scotland (Figure LABEL:fig:sm:areal_data), mirroring the application detailed in the upcoming Section 5.1. This region exhibits a range of Hausdorff distances between zones, from 0.7 to 37.7 kilometers (median: 9.4 km). Data were generated under a Poisson GLMM with a log link function, incorporating a single explanatory variable drawn from a standard normal distribution. The intercept and regression coefficients associated with this explanatory variable were fixed at 4.28 and 0.33, respectively, to mirror the parameters estimated in the real-world application (Section 5.1).
Two random effect structures were considered: the aggregated Gaussian process (AGP) and the Besag–York–Mollié (Besag et al., 1991, BYM), with model parameters corresponding to their fits to the real data. The AGP assumes a point-level GP aggregated at the areal unit level and is closer in structure to the HGP model. We chose AGP’s model parameters to reflect the HGP results from our real-world application (Section 5.1): a point-level marginal standard deviation ($\sigma$) of 0.19, a smoothness parameter ($\nu$) of 1, and a range parameter ($\rho$) of 2.21. On the other hand, the BYM model represents a spatially correlated random effect closer to the structure of the DAGAR model. The BYM model was conveniently parametrized by the marginal precision (Riebler et al., 2016). We set the marginal standard deviation ($\sigma$) to 0.24 and the spatial dependence parameter ($\zeta$) to 0.43, aligning with estimates from the real data analyses in Section 5.1. Section LABEL:sec:sm:sim of the appendix provides further details on the models used to generate data.
We generated 200 datasets from the Poisson GLMM under each of the two distinct random effect structures. For each simulated dataset, we fit two Bayesian Poisson GLMMs, the first one using an homoscedastic HGP while the second one using a DAGAR prior for the random effects. Both specified the fixed effects correctly with the single covariate. The priors for the parameters in the HGP model were set to be those presented in Section 3.2. The priors for the parameters in the DAGAR model were set to the same as those in(Datta et al., 2019). In particular, we set $\pi(\beta_{0})\propto 1$ and $\beta_{1}\sim N(0,1000^{2})$ for the regression parameters; a Gamma prior with shape 2 and rate 1 was placed on the conditional precision ($\phi$), and a uniform prior between 0 and 1 on the spatial dependence parameter ($\psi$). The GoF was evaluated using the criteria in Section 3.2, namely, LOOIC and WAIC, to compare which model fits the data better.
Figure 2: ECDFs displaying the GoF comparison between HGP and DAGAR models across different data generating processes (AGP: solid line, BYM: dashed line) and model comparison criteria (columns of the panel) based on 200 simulated datasets. Shaded region indicates negligible differences. Negative $\Delta_{G}$ indicates better HGP performance.
Figure 2 shows the HGP consistently fitting the data better than the DAGAR model when data were generated from the AGP. The figure displays empirical cumulative distribution functions (ECDFs) of the differences ($\Delta_{G}$) in LOOIC and WAIC (left and right columns of the panel, respectively) between the models. While no formal thresholds exist for significant differences in the model comparison criteria we used (Vranckx et al., 2021), we consider differences exceeding 4 noteworthy. For AGP-generated data (solid line), our simulation study provides strong evidence that the HGP outperforms the DAGAR in GoF. Specifically, $\Delta_{\textrm{LOOIC}}$ was below 10-10 for 80% of simulated datasets, strongly favoring the HGP. In contrast, evidence favoring DAGAR while exceeding ($\Delta_{\textrm{LOOIC}}>4$) was observed in only 1% of the generated data. Interestingly, the HGP remained competitive even when we generated data from the BYM model (a structure closer to the DAGAR). While there was some evidence favoring the DAGAR, $\Delta_{\textrm{LOOIC}}$ was below 4 for more than 50% of the simulations under this scenario. These results highlight the HGP’s robustness as a disease mapping model, demonstrating superior performance when the data aligns with its structure and remains a viable option even when a specialized competitor better represents the underlying process.
In the appendix (Table LABEL:tab:sm:areal), we also evaluated the HGP’s ability to recover true model parameters under correct specification. We assessed estimation accuracy using Mean Absolute Percentage Error (MAPE) and the frequentist coverage percentage (CP) of 95% credible intervals. All model parameters achieved CPs close to the nominal level, with the lowest being 92.5% for $\rho$, which consistently proved the most challenging to estimate. The second most difficult parameter, $\sigma$, exhibited a maximum MAPE of only 7%.
4.2 Data Fusion
Figure 3: Maps utilized in the data fusion simulation study. The left map displays the region used for fitting the models, while the right map shows the data points used to evaluate the models’ predictive accuracy.
To assess the predictive performance of the HGP against standard data fusion models, we conducted a simulation study mirroring usual data fusion scenarios. Within a square region with an area of $51^{2}$, we generated a map comprising a $17\times 17$ grid-boxes of areal units (each with an area of $3^{2}$) and 35 point-referenced data points. Additionally, we created five extra spatial units for each of the following three types of spatial units used exclusively to evaluate predictions: larger polygons (grid boxes with area of $3^{2}$), smaller polygons (grid boxes with area of $1.5^{2}$), and point-referenced, as displayed in Figure 3. We carefully chose the study region and grid box dimensions so each polygon’s area exceeded one, aligning with typical data fusion problems. This design emulates scenarios where areal units, often derived from satellite or computer models, are more abundant than point-referenced data. Hausdorff distances between sample units in this map ranged from 0.4 to 69.8. We simulate the datasets from a Normal GLMM with an identity link function, no covariates, and a spatial random effect. The overall mean of this GLMM was fixed at 6.28 to mirror the application from Section 5.2. In addition, this model introduces the small scale variance (denoted $\tau^{2}$).
[…]
The AGP model assumes:
[…]
where $\{Z(\mathbf{s})\,:\,\mathbf{s}\in D\}$ is a zero-mean, second-order stationary GP with an isotropic Matérn covariance function with smoothness parameter $\nu=1$. SeeMoraga et al. (2017) for a detailed definition of this function. We simulated AGP data with parameters $\sigma^{2}=2.71^{2}$, $\rho=2.29$, and $\tau^{2}=1\times{10}^{-16}$, again mirroring the real data analysis in Section 5.2. In aggregated models, $\tau$ approaches zero as only the average of homoscedastic independent residuals is observed, which converges in probability to zero.
We then generated 200 datasets from the Normal GLMM under each of the two different spatial random effects’ structures detailed next. For each simulated dataset, we fit two Bayesian Normal GLMMs, one using the aforementioned heteroscedastic HGP prior for the random effect and the other using the AGP prior. When fitting the HGP model, we used the priors highlighted in Section 3.2. In addition, we put a penalized complexity prior (Simpson et al., 2017) on $\tau^{-2}$, such that $\mathbbm{P}(\tau>a)=.05$, where $a$ is half the empirical SD observed in the data. Inferences and predictions were based in 2000 MCMC samples with a 1000 warm-up phase per chain with four parallel chains. The initialization of the algorithm is described in Section 3.2.
Inferences for the model induced by the spatial random effect from Equation (9) were performed using the integrated nested Laplace approximation and the stochastic partial differential equation (Lindgren et al., 2011, SPDE) method. The joint distribution induced by this AGP involve approximating stochastic integrals. Practitioners must define a grid with a fixed number of points within each areal unit to approximate these integrals (Wang and Furrer, 2021). The SPDE uses a mesh to discretize the space to achieve such an approximation efficiently. The number of grid points per polygon will depend on the resolution of the mesh. We used two different meshes to show the inferences and, consequently, the predictions may change substantially depending on the accuracy of the approximation for the stochastic integrals in Equation (9). In the appendix , Figure LABEL:fig:sm:meshes-sim visualizes these meshes: a sparse mesh (left column), and a fine-resolution mesh (second column), with approximately 2 and 25 points per polygon, respectively. Henceforth, we will differentiate the AGP models by mesh density, denoting the sparse-mesh model as AGP1 and the fine-mesh model as AGP2. AGP model priors are detailed in Section LABEL:sec:sm:add-sim-fused of the appendix .
We assess the quality of model predictions across different scenarios using four metrics. Firstly, point and interval predictions are derived from the posterior predictive distributions (Equation (8)), using their median and the 2.5% and 97.5% percentiles, respectively. The root mean square error of prediction (RMSP) is used to evaluate point estimates. The remaining three metrics assess interval predictions: the frequentist coverage percentage of prediction intervals, denoted CPP, their width, and the interval score (Gneiting and Raftery, 2007, IS). The latter is defined as
${\rm IS}(y;l,u)=u-l+\frac{2}{\alpha}(l-y)\mathbbm{1}\{y<l\}+\frac{2}{\alpha}(y-u)\mathbbm{1}\{y>u\},$
where $l$ and $u$ are the endpoints of the symmetric $(1-\alpha)\%$ prediction intervals for the target $y$.
Table 1: Out-of-sample prediction for the three models under different data generating models. Spatial unit indicates the geometry of the sample units for which predictions were assessed; for polygons, area is additionally specified within parenthesis. Fit specifies the model fitted to the data. RMSP is the root mean squared error of prediction, CPP is the frequentist coverage percentage of the 95% prediction interval, Width is the prediction interval width, and IS is the interval score. Double headers denote the true data-generating model.
AGP HGP
Spatial unit Fit RMSP CPP Width IS RMSP CPP Width IS
Overall AGP1 1.56 67.8 3.10 17.74 1.81 57.9 2.09 31.37
AGP2 1.39 95.7 5.59 6.60 1.80 88.1 5.53 10.63
HGP 1.65 90.8 6.10 8.65 1.51 95.7 6.15 7.05
Polygons ($3^{2}$) AGP1 0.78 75.8 1.91 5.79 0.89 72.5 1.94 6.96
AGP2 0.60 96.3 2.50 2.87 0.80 89.7 2.68 4.14
HGP 0.83 97.9 3.78 4.00 0.74 96.4 3.00 3.36
Polygons ($1.5^{2}$) AGP1 1.33 75.6 3.56 11.73 0.90 75.6 2.12 6.47
AGP2 1.13 95.8 4.69 5.38 1.09 96.6 4.72 5.32
HGP 1.50 79.2 3.84 9.36 0.85 94.8 3.36 4.01
Points AGP1 2.56 52.2 3.85 35.70 3.64 22.8 2.20 80.70
AGP2 2.43 95.1 9.59 11.54 3.51 77.9 9.19 22.42
HGP 2.62 95.3 10.7 12.57 2.94 96.0 12.09 13.79
Table 1 presents the prediction assessment metrics under both data-generating models at different types of spatial units. Let us first consider the “overall” case, that is, ignoring the types. When the AGP is the true model, the fine mesh AGP model yields the most accurate point predictions (lowest RMSP). However, the HGP model’s performance is only marginally worse, with a RMSP 18% higher. For interval predictions, the HGP model produces prediction intervals (PI) with near-nominal CPP and widths comparable to the AGP2 model, resulting in a competitive interval score (IS). Despite having the narrowest PIs on average, the sparse-mesh AGP model suffers from poor coverage and, consequently, a high IS. As expected, when the true data-generating model is the HGP, the HGP model’s predictive performance surpasses that of its competitors. The point predictions from both AGP models are remarkably similar, with the fine-mesh AGP exhibiting a RMSP only 19% higher than the correctly specified HGP. This suggests the fine-mesh AGP remains competitive in terms of point predictions, despite model misspecification. In terms of interval predictions, while the AGP2 was competitive, its sparse mesh counterpart delivered poor interval predictions. The performance of the competing method varies considerably depending on the mesh resolution, which can significantly impact prediction accuracy. In contrast, our model eliminates the need for arbitrary mesh selection and consistently delivers highly competitive predictions, enhancing prediction accuracy.
Table 1 further reveals demonstrates a consistent decline in prediction accuracy for all models as the area of the spatial units decreased. For AGP-generated data, the AGP2 model’s RMSP for point-referenced data was approximately four times higher than for the polygons with bigger area (i.e., $3^{2}$). Similarly, the HGP model’s RMSP increased by a factor of 3.15 for the same comparison (point-referenced vs largest polygon). For both the HGP and AGP2, CPP remained near the nominal level for all the spatial units. However, HGP’s CPP dropped below nominal level for smaller polygons (area of $1.5^{2}$). This is likely due to variance underestimation: while the AGP-generated data’s variance is inversely proportional to spatial units, the fitted HGP model only distinguish between points and polygons, not their area. Regarding the IS, both our model and the AGP2 model showed IS values roughly three times higher for point-referenced data compared to the largest polygons. This aligns with the simulation parameters. The AGP1 model’s IS performance deteriorated substantially when decreasing the area of the spatial units.
Under the HGP data-generating process, the HGP model consistently outperformed competitors on all assessment metrics except PI width. Although point predictions were broadly similar across models, the HGP provided remarkably superior interval predictions for all spatial unit types, mantaining CPP near the nominal level (see Table 1. In contrast, the AGP1 model exhibited a CPP of only 22.8% for point-referenced data. The poor coverage for both AGP models with point-referenced data is further reflected in their IS, which were 5.9 (AGP1) and 1.6 (AGP2) times higher than the HGP. On the other hand, the AGP2 model remained competitive with the HGP for polygons interval predictions, while the sparse AGP performed poorly.
Table LABEL:tab:sm:fusion in the appendix presents the results for parameter estimation under the correct specification of the the HGP model. We used the same metrics as in Section 4.1 to assess the parameters estimation. The MAPE for $\rho$ was around $10.9\%$ while for $\sigma_{a}$ it was $9.21\%$. Similarly to what is observed with geostatistical models, the small scale standard deviation parameter, namely $\tau$, was the most challenging parameter to be estimated. Unlike the areal simulation, the coverage of some of the variance parameters was consistently below the nominal level of $95\%$. This results are expected, as the HGP inherits the weak identifiability of these parameters from geostatistical models under a normal likelihood (Zhang, 2004). Predictions remain trustworthy despite this issue.
5 Applications
5.1 Areal Data: Respiratory Disease Hospitalization
We first analyze a dataset on the count of hospital admissions due to respiratory disease in 2010 for the 134 intermediate zones (IZ) to the north of the river Clyde in the Great Glasgow and Clyde health board in Scotland. This dataset is publicly available in theCARBayesdata package (Lee, 2022). For each IZ, the expected number of admissions based on the national age- and sex-standardized rates, alongside the percentage of people classified as income deprived in each IZ are also available. The standardized morbidity ratio (SMR), the ratio between the number of observed and expected cases, is often used to control for confounding factors when estimating risks (Greenland, 1984). The goals of the analyses are: (1) to provide smooth estimates of the SMR, adjusted for the percentage of income deprivation, and (2) to contrast the outcomes from the proposed methodology with other established methods.
In this application, $Y(\mathbf{s}_{i})$ is the number of admissions by respiratory disease registered at the $i$-th IZ, denoted $\mathbf{s}_{i}$. In the $i$-th IZ, define $x_{i}$ as the percentage of people who are defined to be income deprived and $E_{i}$ as the expected number of admissions. We analyze the data using the GLMM model from Section 3.1, with a Poisson likelihood, a $\log$ link function, an intercept, a covariate $x_{i}$, and an offset $E_{i}$. We employed the homoscedastic HGP (with a PEC function with smoothness $\nu=0.7$) as the prior for the random effects; the choice of $\nu=0.7$ is justified in Table LABEL:tab:sm:gof_scot of the appendix. The priors for the parameters in the HGP GLMM were set to be those presented in Section 3.2. Inference was performed using four parallel chains with 5000 MCMC samples, thinned by 10, after discarding the initial 1000 iterations as the warm-up period. The initialization of the algorithm is described in Section 3.2. The highest split-$\hat{R}$ was smaller than $1.01$, providing evidence of convergence.
Two competing models, BYM and DAGAR, were considered as the priors for the spatial random effects $\mathbf{z}$ for comparison, as outlined in Sections 4.1 and LABEL:sec:sm:sim, respectively. In the BYM model, parametrized by the marginal precision (Riebler et al., 2016), we followedMorris et al. (2019) and set standard normal priors for the intercept ($\beta_{0}$), regression coefficient ($\beta_{1}$), and the $\mathrm{logit}$ of the mixing parameter $\zeta$. Similarly, a standard normal distribution truncated on positive real numbers was used for $\sigma$. The priors for the DAGAR model are the ones presented in Section 4.1. The initialization of the MCMC algorithm followed the same principle used for the HGP model. In both cases, we used four parallel chains with a warm-up period of 4000 iterations and based the inferences on 20000 MCMC samples, thinned by 20. For these models, the parameters split-$\hat{R}$ provided evidence of convergence. In addition, trace and density plots for posterior samples of all fitted model parameters are provided in Section LABEL:sec:sm:areal_convergence of the appendix.
Table 2: Parameter estimates and model comparison criteria for the three models fitted to the areal dataset. Marginal posterior summary through median and 95% credible intervals.
HGP DAGAR BYM
$\beta_{0}$ -0.21 (-0.268, -0.139) -0.26 (-0.450, -0.137) -0.22 (-0.254, -0.184)
$\beta_{1}$ 0.33 (0.284, 0.368) 0.31 (0.258, 0.370) 0.33 (0.285, 0.366)
$\sigma$ 0.19 (0.155, 0.234) 0.30 (0.218, 0.484) 0.24 (0.176, 0.354)
$\rho$ 2.25 (0.159, 6.948)
$\psi$ 0.43 (0.069, 0.827)
$\xi$ 0.42 (0.100, 0.815)
LOOIC 1081.0 1081.9 1089.3
WAIC 1038.0 1032.4 1039.6
We present the posterior medians and 95% CIs for the parameters from the HGP, DAGAR, and BYM models in Table 2. When considering the coefficient linked to the covariate representing income deprivation, all three models yield similar posterior estimates. The positive posterior estimates the respective parameter ($\beta_{1}$) indicate a positive association between proportion of income deprived people and hospitalization rates for respiratory diseases. Among these models, DAGAR exhibits the broadest credible intervals for the intercept, whereas the marginal posterior distribution for this parameter in the BYM model is highly concentrated around 0.22-0.22. Examining the spatial aspect, the confidence interval for the practical range parameter $\rho$ in the HGP model implies that the distance at which the spatial correlation between random effects from two distinct spatial units decreases to $0.10$ is at most $6.95$ kilometers. Conversely, the 95% CI for the spatial dependence parameter $\psi$ from the DAGAR model spans from $0.07$ to $0.83$. In other words, this credible interval suggests that the clustering effect ranges from almost negligible to very strong. A similar pattern is observed for the $\zeta$ in the BYM model.
Figure 4: Density histogram of the pairwise Hausdorff distances observed in the dataset and. In blue, estimated posterior density of the practical range $\rho$, and, in red, the prior used for $\rho$.
To gain a better understanding of this value in the context of the data, Figure 4 displays a density histogram of all pairwise Hausdorff distances observed in the dataset. The plot also features the prior distribution (in red) and estimated posterior density (in blue) associated with the practical range. Notably, the majority of the posterior distribution mass for $\rho$ falls below the concentration of most pairwise distances, indicating a negligible clustering effect in hospital admissions. An additional analysis of the spatial correlation is presented in Section LABEL:sec:sm:areal_comp in the appendix .
In Table 2, we also present two model comparison criteria described in Section 4 for the models fitted to the data. These results suggest the HGP stands out as the best fit based on the LOOIC, rendering the HGP prior a competitive fit for these data. Despite the modest improvement in model fitting capacity, the HGP possesses an advantage in terms of intuitive interpretability of its spatial dependence, achieved in the last two paragraphs through the practical range parameter $\rho$. Among the specialized models for areal data, the BYM was the least adequate in this application. However, none of the models delivered a remarkably better fit to this dataset. This is further reinforced by looking at the maps of observed and smoothed SMR estimates presented in Figure LABEL:fig:sm:smr from the appendix.
5.2 Data fusion: Air pollution in Ventura and Los Angeles
Figure 5: PM2.5 in Ventura and Los Angeles counties in California, United States. Points represent direct measurements. Tiles are satellite derived estimates.
Our second application analyzes fused fine particulate matter (PM2.5) data in Ventura and Los Angeles counties, California, USA. The dataset combines point-referenced direct measurements (EPA, 2018) from 1999 to present with areal satellite-derived estimates (Van Donkelaar et al., 2015), averaged from 2010 to 2012 (Figure 5). To align temporally, we averaged the direct measurements over the same period. The study region comprises 19 measurement stations and 184 grid boxes (approximately 101.95 km2 each), excluding islands. Our primary goal is to predict PM2.5 concentrations at a finer scale (approximately 45 km2) using both data sources. Secondarily, we aim to estimate average PM2.5 concentration and its spatial dependence across both counties.
Building on previous concepts, we introduce the notation for this application. The response variable is the recorded PM2.5, noted $Y(\mathbf{s}_{i})$ at the spatial unit $\mathbf{s}_{i}$, which is either a point or a grid box. We again resort to the GLMM model from Section 3.1. In this case, there are no covariates, the likelihood is Normal with variance $\tau^{2}$, and the link function is the identity. A heteroscedastic HGP prior is placed on the random effect allowing for differing variances between point-referenced and areal data through the SD function from Equation (2) with $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)\}$. For interpretability, we present results for $\sigma=\exp\{\alpha_{0}\}$ and $\sigma_{a}=\exp\{\alpha_{0}+\alpha_{1}\}$ instead of $\alpha_{0}$ and $\alpha_{1}$. The model was fit with $\nu$ fixed at different values, and we selected the model with best goodness-of-fit ($\nu=0.8$). For more details, see Table LABEL:tab:sm:gof_la in the appendix. The priors for model parameters are as in Section 4.2. The MCMC algorithm initialization is described in Section 3.2. We used four parallel chains, each comprising a warm-up period of 2000 samples and an additional 2000 samples for inference. All model parameters exhibited split-$\hat{R}$ values below, indicating good convergence. This is further supported by the trace plots presented in Section LABEL:sec:sm:df_convergence of the appendix.
For comparison, we also fit an AGP random effect model, with priors for the model parameters as described in Section 4.2. Two meshes of different resolutions were considered; see Figure LABEL:fig:sm:meshes in the appendix. The sparse- and fine-mesh AGP models are denoted AGP1 and AGP2, respectively.
Table 3: Parameter estimates and out-of-sample prediction assessment for the three models fitted to the fused dataset. Marginal posterior summary through median and 95% credible intervals. RMSP is the root mean squared error of prediction; CPP is the frequentist coverage of the 95% prediction interval; Width is the width of the prediction interval; and IS is the interval score.
HGP $\rm AGP_{1}$} $\rm AGP_{2}$}
$\beta$ 5.605 (4.688, 6.449) 6.219 (2.16, 10.10) 6.187 (5.879, 6.478)
$\rho$ 13.826 (7.817, 23.608) 13.163 (5.140, 30.244) 0.632 (0.456, 0.832)
$\tau$ 0.178 (0.073, 0.312) 1.394 (1.242, 1.563) 0.539 (0.394, 0.719)
$\sigma$ 3.849 (2.921, 5.181) 1.703 (0.960, 2.910) 2.399 (2.024, 2.837)
$\sigma_{a}$ 1.241 (1.037, 1.505)
RMSP 1.05 1.45 1.64
Width 3.57 2.53 9.67
CPP 95.5 78.6 95.5
IS 4.80 15.11 13.65
We present the posterior median and 95% CI of the model parameters in Table 3. When looking at the common parameters between the three models (namely, $\beta_{0}$, $\sigma$, $\rho$, and $\tau$), the HGP acts as a compromise between AGP1 and AGP2 while eliminating the arbitrary choice of the level of discretization for approximating numerical integrals. Unlike the aggregated models, our method allows the data to tell us whether the variance of the underlying spatial parameter is smaller or larger than the same quantity for point-referenced data. Interestingly, the estimated SD of the point-referenced data is 3.09 times larger then the areal data (95% CI: 2.43, 4.12). The interpretation of $\rho$, presented in tens of kilometers, in both models is carried out in a similar fashion because the Hausdorff distance between points is equivalent to the distance between points used by the AGP models.
Table 3 also presents the out-of-sample prediction performance of the three models, assessed via 10-fold cross-validation. Using four metrics from our prior simulation study (RMSP, CPP, prediction interval width, IS), we find the HGP offers the most reliable point predictions, evidenced by its lowest RMSP. The HGP also maintains nominal coverage for its prediction intervals. While the AGP2 achieves similar coverage, its prediction intervals are considerably wider, resulting in inferior interval scores. Further analysis by data type is detailed in Table LABEL:tab:sm:out-pred_df of the appendix, revealing the unreliability of competing methods for point-referenced data, where frequentist coverage can be as low as 3%.
The drastic differences in the AGP results for different mesh resolutions are striking. The credible interval width for the intercept varies significantly between the two models. Crucially, the practical range reveals a key divergence: AGP1 indicates moderate spatial dependence, while AGP2 suggests weak dependence, likely explaining the observed overfitting in the latter. This weak dependence may imply AGP2 assigns nearly independent random intercepts to each data point, with variances inversely proportional to the sample unit’s area. AGP mesh resolution is typically practitioner-informed, and results are highly sensitive to this tuning parameter, which lacks a consensus on “optimal resolution” (Fuentes and Raftery, 2005; Liu et al., 2011; Wang and Furrer, 2021).
Figure 6: Change-of-support of PM2.5 to a finer resolution using the sparse and fine AGP and the HGP priors.
Figure 6 displays the change-of-support predictions provided by the three models we fitted to the data. HGP predictions indicate higher PM2.5 concentrations in the southeast (Los Angeles) and provide precise estimates with uncertainty quantification in areas lacking measurement stations, such as the northwestern part of the map. Unlike its counterparts, the HGP’s ability to generate parsimonious predictions without arbitrary discretization, combined with its interpretable parameters, makes it an attractive alternative for spatial data fusion problems.
6 Discussion
The proposed framework unifies the modeling of spatial data, extending geostatistical models to include areal and fused data within the same setting. We established conditions under which the PEC function of the HGP is positive definite, ensuring its validity. As a prior for spatial random effects in a GLMM, the HGP proves to be competitive to specialized models for diverse spatial data types. Our simulation studies indicate that the HGP is particularly advantageous for areal data with units of varying sizes and shapes, as confirmed by our disease mapping application. Unlike its specialized counterparts, the HGP also provides a directly interpretable spatial parameter, offering insights into the practical range of spatial dependence. In data fusion contexts, the HGP delivers more reliable interval predictions compared to the sparse AGP, and it provided similar results to the fine AGP in parameter estimation and spatial range. Additionally, in the air pollution data application, the proposed model consistently outperformed the AGP regardless of the resolution of its mesh. This success is even more remarkable considering the HGP bypasses the subjective step of defining a grid or mesh, a crucial but problematic requirement for the AGP. In addition, independent of the spatial data type, it continues to have direct and interpretable parameters.
Our proposed methodology shares the “big nn” computational complexity challenge inherent in traditional geostatistical approaches. Consequently, the computational cost of evaluating the finite-dimensional likelihood induced by the HGP for a set of $n$ areal units scales as $\mathcal{O}(n^{3})$. Proving the validity of the PEC function under the Hausdorff distance on Euclidean space was a demanding theoretical task. Relaxing the isotropy assumption for this function presents further theoretical complexities. Moreover, to apply the HGP to data measured on manifolds (like the sphere), new theoretical results are needed to guarantee the positive definiteness of a correlation function within that specific geometric context.
The HGP framework offers numerous opportunities for development. Further researching its theoretical properties will strengthen the foundation for its applications. A crucial focus should be developing correlation functions designed specifically for the HGP, particularly those suited to scenarios where spatial data is observed on a sphere. Approximations inspired by methods like the Nearest Neighbor Gaussian Process (Datta et al., 2016; Quiroz et al., 2023) or reduced rank approximations (Solin and Särkkä, 2020) could significantly broaden the applicability of the HGP to large-scale datasets. These advancements hold the potential to establish the HGP as an even more powerful and versatile tool for spatial data analysis.

### v3

@@
-Statistical Inferences and Predictions for Areal Data and Spatial Data Fusion with Hausdorff–Gaussian Processes
+Belo Horizonte, MG, Brazil
 Abstract
-Accurate modeling of spatial dependence is pivotal in analyzing spatial data, influencing parameter estimation and out-of-sample predictions.
-The spatial structure and geometry of the data significantly impact valid statistical inference.
+Accurate modeling of spatial dependence is pivotal in analyzing spatial data, influencing parameter estimation and predictions.
+The spatial structure of the data significantly impacts valid statistical inference.
 Existing models for areal data often rely on adjacency matrices, struggling to differentiate between polygons of varying sizes and shapes.
-Conversely, data fusion models, while effective, rely on computationally intensive numerical integrals, presenting challenges for moderately large datasets.
-In response to these issues, we propose the Hausdorff-Gaussian process (HGP), a versatile model class utilizing the Hausdorff distance to capture spatial dependence in both point and areal data.
-We introduce a valid correlation function within the HGP framework, accommodating diverse modeling techniques, including geostatistical and areal models.
-Integration into generalized linear mixed-effects models enhances its applicability, particularly in addressing change of support and data fusion challenges.
+Conversely, data fusion models rely on computationally intensive numerical integrals, presenting challenges for moderately large datasets.
+In response to these issues, we propose the Hausdorff-Gaussian process (HGP), a versatile model utilizing the Hausdorff distance to capture spatial dependence in both point and areal data.
+Integration into generalized linear mixed-effects models enhances its applicability, particularly in addressing data fusion challenges.
 We validate our approach through a comprehensive simulation study and application to two real-world scenarios: one involving areal data and another demonstrating its effectiveness in data fusion.
@@
 In summary, the HGP offers a flexible and robust solution for modeling spatial data of various types and shapes, with potential applications spanning fields such as public health and climate science.
-Keywords: Areal data, CAR models, Geostatistics, Spatial modeling, Spatial prediction
+Keywords: CAR models; Geostatistics; Spatial data fusion; Spatial modeling; Spatial prediction
 1 Introduction
@@
 Valid statistical inference are deeply influenced by the spatial structure and geometry of the data.
-Point-referenced data are typically analyzed using geostatistics methods (e.g., Cressie, 1993).
-For areal data, the conditional autoregressive model (CAR) (Besag, 1974) and its variants (Besag et al., 1991; Rodrigues and Assunçao, 2012; Datta et al., 2019; Cruz-Reyes et al., 2023) are often preferred.
+Point-referenced data are typically analyzed using geostatistics methods (e.g., Cressie,, 1993).
+For areal data, the conditional autoregressive model (Besag,, 1974, CAR) and its variants (Besag et al.,, 1991; Rodrigues and Assunçao,, 2012; Datta et al.,, 2019; Cruz-Reyes et al.,, 2023) are often preferred.
 The nature of the observed spatial data significantly influences the choice of statistical methodology.
 This becomes increasingly critical in complex inferential problems, such as predicting a variable in a target map using data from a source map.
-This issue, known as change of support (Cressie, 1993, pg.
-284–289), requires tailored methods depending on the spatial units involved (Gelfand et al., 2001).
-Spatial data from diverse sources often differ in spatial units or supports, resulting in spatial misalignment (Gotway and Young, 2002).
-Data fusion is necessitated when a spatial variable is provided by multiple, potentially misaligned, sources (Moraga et al., 2017).
+This issue, known as change of support (Cressie,, 1993, pg.
+284–289), requires tailored methods depending on the spatial units involved (Gelfand et al.,, 2001).
+Spatial data from diverse sources often differ in spatial units or supports, resulting in spatial misalignment (Gotway and Young,, 2002).
+Data fusion is necessitated when a spatial variable is provided by multiple, potentially misaligned, sources (Moraga et al.,, 2017).
 Both traditional areal models and data fusion models for spatial data exhibit inherent limitations.
-Traditional areal models face challenges with predictions for overlapping spatial units (Bakar and Jin, 2020), do not account for region size or shape of the areal units and lack interpretable spatial dependence parameters (Assunção and Krainski, 2009; Cruz-Reyes et al., 2023).
-An exception of the last limitation is the directed acyclic graph auto-regressive (DAGAR) model, which offers an interpretable parameter (Datta et al., 2019).
-Data fusion models, which view areal data as averages of a continuous process, may be unsuitable for discrete outcomes (Gelfand et al., 2001).
-Additionally, their reliance on arbitrary discretization for likelihood evaluation lacks consensus on optimal resolution (Fuentes and Raftery, 2005; Liu et al., 2011; Wang and Furrer, 2021) and can introduce unquantifiable biases (Gonçalves and Gamerman, 2018).
-We propose a unified approach for analyzing spatial data encompassing areal, point-referenced, and mixed spatial supports.
-This methodology combines Gaussian processes (GPs) with the Hausdorff distance (Shephard and Webster, 1965).
-The Hausdorff distance, defined on a reference metric space, effectively measures between the distance between two sets while accounting for their shape and size.
+Traditional areal models face challenges with predictions for overlapping spatial units (Bakar and Jin,, 2020), do not account for region size or shape of the areal units and lack interpretable spatial dependence parameters (Assunção and Krainski,, 2009; Cruz-Reyes et al.,, 2023).
+An exception of the last limitation is the directed acyclic graph auto-regressive (DAGAR) model, which offers an interpretable parameter (Datta et al.,, 2019).
+Data fusion models, which view areal data as averages of a continuous process, may be unsuitable for discrete outcomes (Gelfand et al.,, 2001).
+Additionally, their reliance on arbitrary discretization for likelihood evaluation lacks consensus on optimal resolution (Fuentes and Raftery,, 2005; Liu et al.,, 2011; Wang and Furrer,, 2021) and can introduce unquantifiable biases (Gonçalves and Gamerman,, 2018).
+We introduce a unified and flexible approach for analyzing spatial data of varying types, encompassing areal, point-referenced, and mixed spatial supports.
+Our key objective is to develop a model capable of seamlessly handling these diverse data formats within a single framework.
+This is achieved by combining Gaussian processes (GPs) with the Hausdorff distance (Arutyunov et al.,, 2016).
+The Hausdorff distance, defined on a reference metric space, effectively measures the distance between two sets while accounting for their shape.
 For point pairs, it reduces to the distance metric associated to the reference metric space.
-GPs, recognized for their flexibility, are widely utilized in spatial data modeling, notably as priors in generalized mixed effects models (Diggle et al., 1998), for spatially varying regression coefficients (Gelfand et al., 2003), and in addressing spatial misalignment issues (Banerjee and Gelfand, 2002).
-The Hausdorff–Gaussian process (HGP) that we propose is a GP with a correlation function defined on the Hausdorff distance between sample units, allowing integrating their shape and size into the analysis.
-With some ensurance of positiveness definiteness of the resulting covariance matrix, the HGP with a powered exponential correlation can be used in place of GPs to adeptly model areal, point-referenced, and mixed spatial data types, offering a new avenue in analyzing spatial and spatiotemporal data with mixed supports.
+GPs, recognized for their flexibility, are widely utilized in spatial data modeling, notably as priors in generalized mixed effects models (Diggle et al.,, 1998), in spatially varying regression coefficients (Gelfand et al.,, 2003), and in addressing spatial misalignment issues (Banerjee and Gelfand,, 2002).
+We propose the Hausdorff-Gaussian process (HGP), a GP with a correlation function defined on the Hausdorff distance between spatial units.
+This allows us to directly incorporate shape and size information into the analysis.
+A crucial aspect of our work is to assess the competitiveness of the HGP, particularly when compared to specialized models designed for areal and mixed spatial data.
+We investigate whether the HGP, using a powered exponential correlation, can effectively model areal and mixed spatial data, thereby offering a novel and versatile tool for analyzing spatial and spatiotemporal data with mixed supports.
 This paper is structured as follows.
@@
 2.1 Hausdorff Distance
-The Hausdorff distance is a metric measuring the distance between two sets of points.
+The Hausdorff distance is a metric measuring the distance between two sets.
 Specifically, it is the greatest distance from a point in one set to the closest point in the other set.
@@
 For a single point $x$ and a set $A$, define $d(x,A)=\inf_{a\in A}d(x,a)$ as the closest distance from $x$ to any element in $A$.
-The directed Hausdorff distance from a set $A$ to a set $B$ is $\vec{h}(A,B)=\sup_{a\in A}d(a,B).$ This measure is not necessarily symmetric.
-The symmetric Hausdorff distance is defined as the greater of the two directed Hausdorff distances: $h(A,B)=\max\left\{\vec{h}(A,B),\vec{h}(B,A)\right\}.$ Note that if $A$ and $B$ are both singletons, then $h(A,B)=d(A,B)$.
-In spatial analysis, the Hausdorff distance provides a robust and suitable metric over non-empty, bounded, and compact subsets of $D$, denoted as $\mathcal{B}(D)$ (Sendov, 2004).
+The directed Hausdorff distance from a set $A$ to a set $B$ is
+$\vec{h}(A,B)=\sup_{a\in A}d(a,B).$
+This measure is not necessarily symmetric. The symmetric Hausdorff distance is defined as the greater of the two directed Hausdorff distances:
+$h(A,B)=\max\left\{\vec{h}(A,B),\vec{h}(B,A)\right\}.$
+Note that if $A$ and $B$ are both singletons, then $h(A,B)=d(A,B)$. Interestingly, if $A$ and $B$ are compact, there exist $a\in A$ and $b\in B$ such that $h(A,B)=d(a,b)$ (Arutyunov et al.,, 2016).
+In spatial analysis, the Hausdorff distance provides a robust and suitable metric over non-empty compact subsets of $D$, denoted as $\mathcal{B}(D)$ (Molchanov,, 2005, Appendix C).
 While not functioning as a conventional distance in unbounded sets, it is highly effective in bounded contexts, which are prevalent in spatial statistics.
 Typically, the study region $D$ in spatial statistics is a bounded subset of $\mathbb{R}^{2}$, where the Hausdorff distance excels in quantifying distances between spatial units by capturing both their shapes and sizes.
+The metric space induced by the tuple $(\mathcal{B}(D),h)$ inherits properties like compactness and completeness from $(D,d)$.
+Further details regarding the topology induced by the Hausdorff distance can be found in Appendix C inMolchanov, (2005).
 AB
@@
 In each sub-figure, set $A$ is represented as a circle with a radius of 1.
-Figure 1(a) introduces a set $B$, a circle tangent to $A$.
-Figure 1(b) examines the distance between two sets having a non-empty intersection.
-Figure 1(c) features $B$ as an interior point of $A$.
+Figure 1a introduces a set $B$, a circle tangent to $A$.
+Figure 1b examines the distance between two sets having a non-empty intersection.
+Figure 1c features $B$ as an interior point of $A$.
 The dashed lines in the figures delineate the Hausdorff distances between the sets, highlighting the metric’s ability to consider the shape, size, and orientation of the objects.
@@
 $\log v(\mathbf{s})=\alpha_{0}+\alpha_{1}w(\mathbf{s}).$ (2)
-Equation (2) allows a distinction between point-referenced and areal data, a need in data fusion applications (Wang and Furrer, 2021).
+Equation (2) allows a distinction between point-referenced and areal data, a need in data fusion applications (Wang and Furrer,, 2021).
 With $\mathcal{A}(\cdot)$ being the function that returns the area of a spatial object.
 A covariate $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$ enables different variances between these data types.
-Certainly, care is necessary in specifying $v(\mathbf{s})$ to ensure a valid GP (Palacios and Steel, 2006).
+Certainly, care is necessary in specifying $v(\mathbf{s})$ to ensure a valid GP (Palacios and Steel,, 2006).
 2.3 Correlation Function
@@
 Therefore, unless explicitly mentioned, all subsequent references to correlation functions will assume they are isotropic.
-A valid correlation function must be positive definite and meet two key criteria: $r(0)=1$ and $\lim_{h\to\infty}r(h)=0$ (Cressie, 1993, Ch.
+A valid correlation function must be positive definite and meet two key criteria: $r(0)=1$ and $\lim_{h\to\infty}r(h)=0$ (Cressie,, 1993, Ch.
 2.3.2), where $h$ is the Hausdorff distance between a pair $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{B}(D)$.
 The latter ensures that spatial correlation diminishes with increasing distance.
-It is noteworthy that while most geostatistical correlation functions are constructed for Euclidean spaces (Cressie, 1993, Ch.
-2.5.1), their positive definiteness may not extend to non-Euclidean spaces (Curriero, 2006; Li et al., 2023).
-This raises the question of whether existing frameworks can be adapted or if entirely new formulations are needed to maintain the mathematical rigor in non-Euclidean settings.
+It is noteworthy that while most geostatistical correlation functions are constructed for Euclidean spaces (Cressie,, 1993, Ch.
+2.5.1), their positive definiteness may not extend to non-Euclidean spaces (Curriero,, 2006; Li et al.,, 2023).
 For practical purposes, we propose to use a two-parameter powered exponential correlation (PEC) function
-$r(h;\varphi,\nu)=\exp\left\{-\frac{h^{\nu}}{\varphi^{\nu}}\right\},$
+$r(h;\varphi,\nu)=\exp\left\{-\left(\frac{h}{\varphi}\right)^{\nu}\right\},$
 where $\nu$ controls the smoothness and $\varphi$ governs the range of the spatial dependence.
@@
 Consequently, $\rho$ is the distance at which the spatial correlation between a pair of spatial units drops to $0.10$, commonly referred to as the practical range.
-The PEC function is appealing because the following Proposition 2.1 guarantees its positive definiteness for $\nu\in(1/2,1)$.
-The proof of the proposition is based on probabilistic embedding of Hausdorff metrics (Backurs and Sidiropoulos, 2016).
-Proposition 2.1.
-Let $h=h(\mathbf{s},\mathbf{s}^{\prime})$ be the Hausdorff distance between two spatial units, denoted $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{B}(D)$, where $D\subset\rm I\!R^{2}$.
-The powered exponential correlation function $\exp\{-h^{\nu}/\varphi^{\nu}\}$ is positive definite for $\nu\in(1/2,1)$.
-In addition to the theoretical result, an empirical assessment was used to provide evidence for the positive definiteness of this function for an expanded range of $\nu\in(0,1)$.
-The positive definiteness is satisfied for bigger $\nu$ values when the dependence range $\rho$ is smaller.
-See details in Section LABEL:sec:sm:emp of the appendix.
+The PEC function is appealing because, for $\nu\in(0,1]$, it is positive definite as long as the distance function is conditionally negative definite (Schoenberg,, 1938).
+To the best of our knowledge, there exist no theoretical results on positive definite functions for the Hausdorff distance.
+Therefore, we recommend an empirical assessment of the PEC function positive definiteness before analyzing the data.
+To this end, we empirically assess the positive definiteness of the PEC function for $\nu\in(0,1]$.
+We first calculate Hausdorff distance matrices for the areal and fused data to be analyzed in Sections 5.1 and 5.2, respectively.
+Evaluating the PEC function across these matrices with $\nu\in[0.01,1]$ and $\rho\in[d_{\min},d_{\max}]$ (where $d_{\min}$ and $d_{\max}$ are the dataset’s minimum and maximum distances) yields corresponding correlation matrices.
+We then compute each matrix’s smallest eigenvalue.
+A negative eigenvalue suggests the corresponding PEC function is not positive definite.
+However, even valid covariance functions can be ill-conditioned or nearly rank-deficient at dense locations or when there exists a severely strong spatial dependence (Banerjee et al.,, 2012).
+Figure 2: Smallest eigenvalue of the powered exponential correlation matrix under different combinations of smoothness ($\nu$) and practical range ($\rho$) for areal and fused data.
+The transparency indicates regions where $\lambda_{1}<0$.
+Figure 2 displays the smallest eigenvalues ($\lambda_{1}$) of the PEC-induced correlation matrices for both datasets.
+The fused data exhibits positive eigenvalues for all $(\nu,\rho)$ pairs.
+Conversely, the areal data reveals some small negative eigenvalues (range 0.0099-0.0099 to 0.00006-0.00006) under severely strong spatial dependence and at $\nu\approx 1$, the border of the parametric space.
+Therefore, our empirical study suggests the positive definiteness of the PEC function for $\nu\in(0,1]$.
 3 HGP in Spatial GLMM
@@
 For fixed effects regression coefficients ($\bm{\beta}$), we employ uncorrelated zero-mean Normal priors with marginal variances $10^{2}$.
-In the homoscedastic case, on $\sigma$ we place a half-t prior with $3$ degrees of freedom, denoted $t_{+}(3)$ (Gelman, 2006).
+In the homoscedastic case, on $\sigma$ we place a half-t prior with $3$ degrees of freedom, denoted $t_{+}(3)$ (Gelman,, 2006).
 For the hereroscedastic model (2), we set independent standard normal priors for $\alpha_{0}$ and $\alpha_{1}$.
 The PEC dependence is encoded through $\bm{\delta}=\{\rho,\nu\}$.
-In geostatistics, smoothness parameters for correlation functions are typically fixed, and their estimation properties are rarely reported (Zhang, 2004).
+In geostatistics, smoothness parameters for correlation functions are typically fixed, and their estimation properties are rarely reported (Zhang,, 2004).
 Due to this and the scope of this paper, we specify separate models for different values of $\nu$ ensuring a positive definite PEC and select the best fit for inference.
@@
 The estimation of the spatial GLMM parameters is obtained from the posterior distribution defined in Equation (7).
-As the posterior distribution is intractable, we use the No-U-Turn (Hoffman and Gelman, 2014) Markov Chain Monte Carlo (MCMC) sampler available in Stan (Team, 2023) to draw samples from it.
+As the posterior distribution is intractable, we use the No-U-Turn (Homan and Gelman,, 2014) Markov Chain Monte Carlo (MCMC) sampler available in Stan (Team,, 2023) to draw samples from it.
 The No-U-Turn algorithm samples all the model parameters jointly and efficiently exploits the parameter space using automatic differentiation.
@@
 The number of samples and the warm-up period of the MCMC algorithm are application dependent.
-We assess the convergence of the chains using the split-$\hat{R}$ diagnostic (Vehtari et al., 2021).
+We assess the convergence of the chains using the split-$\hat{R}$ diagnostic (Vehtari et al.,, 2021).
 Finally, the parameters’ point estimates and 95% credible intervals (CI) are obtained as, respectively, the median and percentiles ($2.5$ and $97.5$, unless otherwise stated) of the marginal MCMC samples.
 Model Comparison
-The assessment of goodness-of-fit (GoF) in the simulation study and applications involved two model comparison criteria: the leave-one-out information criterion (Vehtari et al., 2017, LOOIC), and the widely applicable information criterion (Watanabe, 2010, WAIC).
+The assessment of goodness-of-fit (GoF) in the simulation study and applications involved two model comparison criteria: the leave-one-out information criterion (Vehtari et al.,, 2017, LOOIC), and the widely applicable information criterion (Watanabe and Opper,, 2010, WAIC).
 For both of them, lower values indicate a better fit.
-We did not use the deviance information criterion due to its assumption of a small number of parameters relative to the sample size (McElreath, 2016, Ch.
+We did not use the deviance information criterion due to its assumption of a small number of parameters relative to the sample size (McElreath,, 2016, Ch.
 7.4), which would require computationally intensive numerical integration of random effects in some of our models.
@@
 Predictions
-Predicting spatial phenomena accurately at unobserved locations is crucial in many applications, especially in change of support settings (Gelfand et al., 2001).
+Predicting spatial phenomena accurately at unobserved locations is crucial in many applications, especially in change of support settings (Gelfand et al.,, 2001).
 The HGP prior for spatial random effects in GLMMs offers a powerful solution.
@@
 where $\mathbf{z}^{\ast}={(z(\mathbf{s}^{\ast}_{1}),\ldots,z(\mathbf{s}^{\ast}_{m}))}^{\top}$ is a realization of the spatial random effect at the new locations, and $\bm{\theta}$ represents the model parameters.
-Details on sampling from this distribution are provided in Section LABEL:sec:sm:hgp-pred of the appendix.
+Details on sampling from this distribution are provided in Section S.1 of the Appendix Material.
 4 Simulation Studies
 4.1 Areal Data
-We conducted a simulation study to assess the performance of the HGP relative to the DAGAR model, a state-of-the-art model designed explicitly for areal data (Datta et al., 2019), in modeling data generated from other models.
-We grounded our simulations in a real-world context by utilizing the 134 Intermediate Zones (IZ) north of the River Clyde within the Greater Glasgow in Scotland (Figure LABEL:fig:sm:areal_data), mirroring the application detailed in the upcoming Section 5.1.
+We conducted a simulation study to assess the performance of the HGP relative to the DAGAR model, a state-of-the-art model designed explicitly for areal data (Datta et al.,, 2019), in modeling data generated from other models.
+We grounded our simulations in a real-world context by utilizing the 134 Intermediate Zones (IZ) north of the River Clyde within the Greater Glasgow in Scotland (Figure S.1), mirroring the application detailed in the upcoming Section 5.1.
 This region exhibits a range of Hausdorff distances between zones, from 0.7 to 37.7 kilometers (median: 9.4 km).
@@
 The intercept and regression coefficients associated with this explanatory variable were fixed at 4.28 and 0.33, respectively, to mirror the parameters estimated in the real-world application (Section 5.1).
-Two random effect structures were considered: the aggregated Gaussian process (AGP) and the Besag–York–Mollié (Besag et al., 1991, BYM), with model parameters corresponding to their fits to the real data.
+Two random effect structures were considered: the aggregated Gaussian process (AGP) and the Besag–York–Mollié (Besag et al.,, 1991, BYM), with model parameters corresponding to their fits to the real data.
 The AGP assumes a point-level GP aggregated at the areal unit level and is closer in structure to the HGP model.
 We chose AGP’s model parameters to reflect the HGP results from our real-world application (Section 5.1): a point-level marginal standard deviation ($\sigma$) of 0.19, a smoothness parameter ($\nu$) of 1, and a range parameter ($\rho$) of 2.21.
-On the other hand, the BYM model represents a spatially correlated random effect closer to the structure of the DAGAR model.
-The BYM model was conveniently parametrized by the marginal precision (Riebler et al., 2016).
-We set the marginal standard deviation ($\sigma$) to 0.24 and the spatial dependence parameter ($\zeta$) to 0.43, aligning with estimates from the real data analyses in Section 5.1.
-Section LABEL:sec:sm:sim of the appendix provides further details on the models used to generate data.
+The BYM model allows for spatially correlated random effects with a structure closer to that of the DAGAR model.
+We set its marginal standard deviation (Riebler et al.,, 2016), denoted $\sigma$, to 0.24 and the spatial dependence parameter ($\xi$) to 0.43, aligning with estimates from the real data analyses in Section 5.1.
+Section S.2.1 of the Appendix Material provides further details on the models used to generate data.
 We generated 200 datasets from the Poisson GLMM under each of the two distinct random effect structures.
@@
 The priors for the parameters in the HGP model were set to be those presented in Section 3.2.
-The priors for the parameters in the DAGAR model were set to the same as those in(Datta et al., 2019).
+The priors for the parameters in the DAGAR model were set to the same as those in(Datta et al.,, 2019).
 In particular, we set $\pi(\beta_{0})\propto 1$ and $\beta_{1}\sim N(0,1000^{2})$ for the regression parameters; a Gamma prior with shape 2 and rate 1 was placed on the conditional precision ($\phi$), and a uniform prior between 0 and 1 on the spatial dependence parameter ($\psi$).
@@
 The figure displays empirical cumulative distribution functions (ECDFs) of the differences ($\Delta_{G}$) in LOOIC and WAIC (left and right columns of the panel, respectively) between the models.
-While no formal thresholds exist for significant differences in the model comparison criteria we used (Vranckx et al., 2021), we consider differences exceeding 4 noteworthy.
+While no formal thresholds exist for significant differences in the model comparison criteria we used, we consider differences exceeding 5 to be noteworthy, following a similar practice used with the DIC (Vranckx et al.,, 2021).
 For AGP-generated data (solid line), our simulation study provides strong evidence that the HGP outperforms the DAGAR in GoF.
 Specifically, $\Delta_{\textrm{LOOIC}}$ was below 10-10 for 80% of simulated datasets, strongly favoring the HGP.
-In contrast, evidence favoring DAGAR while exceeding ($\Delta_{\textrm{LOOIC}}>4$) was observed in only 1% of the generated data.
+In contrast, evidence favoring DAGAR ($\Delta_{\textrm{LOOIC}}>4$) was observed in only 1% of the generated data.
 Interestingly, the HGP remained competitive even when we generated data from the BYM model (a structure closer to the DAGAR).
@@
 These results highlight the HGP’s robustness as a disease mapping model, demonstrating superior performance when the data aligns with its structure and remains a viable option even when a specialized competitor better represents the underlying process.
-In the appendix (Table LABEL:tab:sm:areal), we also evaluated the HGP’s ability to recover true model parameters under correct specification.
+In the Appendix Material (Table S.1), we also evaluated the HGP’s ability to recover true model parameters under correct specification.
 We assessed estimation accuracy using Mean Absolute Percentage Error (MAPE) and the frequentist coverage percentage (CP) of 95% credible intervals.
 All model parameters achieved CPs close to the nominal level, with the lowest being 92.5% for $\rho$, which consistently proved the most challenging to estimate.
-The second most difficult parameter, $\sigma$, exhibited a maximum MAPE of only 7%.
+The marginal standard deviation ($\sigma$) exhibited a maximum MAPE of only 7%.
 4.2 Data Fusion
@@
 Within a square region with an area of $51^{2}$, we generated a map comprising a $17\times 17$ grid-boxes of areal units (each with an area of $3^{2}$) and 35 point-referenced data points.
-Additionally, we created five extra spatial units for each of the following three types of spatial units used exclusively to evaluate predictions: larger polygons (grid boxes with area of $3^{2}$), smaller polygons (grid boxes with area of $1.5^{2}$), and point-referenced, as displayed in Figure 3.
+Additionally, we created five extra spatial units for each of the following three types exclusively to evaluate predictions: larger polygons (grid boxes with area of $3^{2}$), smaller polygons (grid boxes with area of $1.5^{2}$), and point-referenced, as displayed in Figure 4.
 We carefully chose the study region and grid box dimensions so each polygon’s area exceeded one, aligning with typical data fusion problems.
@@
 In addition, this model introduces the small scale variance (denoted $\tau^{2}$).
-[…]
+This section’s spatial GLMMs employ two distinct spatial random effect structures: a heteroscedastic HGP and an AGP.
+The HGP’s SD function $v(\mathbf{s})$ is defined as $\exp\{\alpha_{0}+\alpha_{1}\mathbbm{1}(\lvert\mathbf{s}\rvert>0)\}$, allowing for distinct variances between point-referenced and areal data.
+We selected HGP parameters ($\tau^{2}=0.17^{2}$, $\sigma=4.4416$, $\sigma_{a}=1.3114$, $\rho=28.45$, and $\nu=0.8$) to closely resemble the data analysis presented in Section 5.2.
+Note that the spatial variance for point and areal units ($\sigma$ and $\sigma_{a}$, respectively) is a one-to-one transformation of $(\alpha_{0},\alpha_{1})$$^{\top}$.
+We set the dependence parameter $\rho$ to 48.9.
+This value differs from our estimate in the real data due to a necessary rescaling based on the map used in this simulation study.
+For further details, refer to Section S.3 in the Appendix Material.
 The AGP model assumes:
@@
 where $\{Z(\mathbf{s})\,:\,\mathbf{s}\in D\}$ is a zero-mean, second-order stationary GP with an isotropic Matérn covariance function with smoothness parameter $\nu=1$.
-SeeMoraga et al.
-(2017) for a detailed definition of this function.
+SeeMoraga et al., (2017) for a detailed definition of this function.
 We simulated AGP data with parameters $\sigma^{2}=2.71^{2}$, $\rho=2.29$, and $\tau^{2}=1\times{10}^{-16}$, again mirroring the real data analysis in Section 5.2.
@@
 When fitting the HGP model, we used the priors highlighted in Section 3.2.
-In addition, we put a penalized complexity prior (Simpson et al., 2017) on $\tau^{-2}$, such that $\mathbbm{P}(\tau>a)=.05$, where $a$ is half the empirical SD observed in the data.
+In addition, we put a penalized complexity prior (Simpson et al.,, 2017) on $\tau^{-2}$, such that $\mathbbm{P}(\tau>a)=.05$, where $a$ is half the empirical SD observed in the data.
 Inferences and predictions were based in 2000 MCMC samples with a 1000 warm-up phase per chain with four parallel chains.
 The initialization of the algorithm is described in Section 3.2.
-Inferences for the model induced by the spatial random effect from Equation (9) were performed using the integrated nested Laplace approximation and the stochastic partial differential equation (Lindgren et al., 2011, SPDE) method.
+Inferences for the model induced by the spatial random effect from Equation (9) were performed using the integrated nested Laplace approximation and the stochastic partial differential equation (Lindgren et al.,, 2011, SPDE) method.
 The joint distribution induced by this AGP involve approximating stochastic integrals.
-Practitioners must define a grid with a fixed number of points within each areal unit to approximate these integrals (Wang and Furrer, 2021).
+Practitioners must define a grid with a fixed number of points within each areal unit to approximate these integrals (Wang and Furrer,, 2021).
 The SPDE uses a mesh to discretize the space to achieve such an approximation efficiently.
-The number of grid points per polygon will depend on the resolution of the mesh.
+The number of grid boxes per polygon will depend on the resolution of the mesh.
 We used two different meshes to show the inferences and, consequently, the predictions may change substantially depending on the accuracy of the approximation for the stochastic integrals in Equation (9).
-In the appendix , Figure LABEL:fig:sm:meshes-sim visualizes these meshes: a sparse mesh (left column), and a fine-resolution mesh (second column), with approximately 2 and 25 points per polygon, respectively.
+In the Appendix Material, Figure S.2 visualizes these meshes: a sparse mesh (left column), and a fine-resolution mesh (second column), with approximately 2 and 25 points per polygon, respectively.
 Henceforth, we will differentiate the AGP models by mesh density, denoting the sparse-mesh model as AGP1 and the fine-mesh model as AGP2.
-AGP model priors are detailed in Section LABEL:sec:sm:add-sim-fused of the appendix .
+Their priors are detailed in Section S.4 of the Appendix Material.
 We assess the quality of model predictions across different scenarios using four metrics.
@@
 The root mean square error of prediction (RMSP) is used to evaluate point estimates.
-The remaining three metrics assess interval predictions: the frequentist coverage percentage of prediction intervals, denoted CPP, their width, and the interval score (Gneiting and Raftery, 2007, IS).
+The remaining three metrics assess interval predictions: the frequentist coverage percentage of prediction intervals, denoted CPP, their width, and the interval score (Gneiting and Raftery,, 2007, IS).
 The latter is defined as
@@
 Table 1: Out-of-sample prediction for the three models under different data generating models.
-Spatial unit indicates the geometry of the sample units for which predictions were assessed; for polygons, area is additionally specified within parenthesis.
+“Sp.
+unit” indicates the geometry of the sample units for which predictions were assessed; for polygons (denoted Poly), area is additionally specified within parenthesis.
 Fit specifies the model fitted to the data.
@@
 AGP HGP
-Spatial unit Fit RMSP CPP Width IS RMSP CPP Width IS
+Sp. unit Fit RMSP CPP Width IS RMSP CPP Width IS
 Overall AGP1 1.56 67.8 3.10 17.74 1.81 57.9 2.09 31.37
@@
 HGP 1.65 90.8 6.10 8.65 1.51 95.7 6.15 7.05
-Polygons ($3^{2}$) AGP1 0.78 75.8 1.91 5.79 0.89 72.5 1.94 6.96
+Poly ($3^{2}$) AGP1 0.78 75.8 1.91 5.79 0.89 72.5 1.94 6.96
 AGP2 0.60 96.3 2.50 2.87 0.80 89.7 2.68 4.14
 HGP 0.83 97.9 3.78 4.00 0.74 96.4 3.00 3.36
-Polygons ($1.5^{2}$) AGP1 1.33 75.6 3.56 11.73 0.90 75.6 2.12 6.47
+Poly ($1.5^{2}$) AGP1 1.33 75.6 3.56 11.73 0.90 75.6 2.12 6.47
 AGP2 1.13 95.8 4.69 5.38 1.09 96.6 4.72 5.32
@@
 In contrast, our model eliminates the need for arbitrary mesh selection and consistently delivers highly competitive predictions, enhancing prediction accuracy.
-Table 1 further reveals demonstrates a consistent decline in prediction accuracy for all models as the area of the spatial units decreased.
+Table 1 further reveals a consistent decline in prediction accuracy for all models as the area of the spatial units decreased.
 For AGP-generated data, the AGP2 model’s RMSP for point-referenced data was approximately four times higher than for the polygons with bigger area (i.e., $3^{2}$).
@@
 Under the HGP data-generating process, the HGP model consistently outperformed competitors on all assessment metrics except PI width.
-Although point predictions were broadly similar across models, the HGP provided remarkably superior interval predictions for all spatial unit types, mantaining CPP near the nominal level (see Table 1.
+Although point predictions were broadly similar across models, the HGP provided remarkably superior interval predictions for all spatial unit types, mantaining CPP near the nominal level (see Table 1).
 In contrast, the AGP1 model exhibited a CPP of only 22.8% for point-referenced data.
@@
 On the other hand, the AGP2 model remained competitive with the HGP for polygons interval predictions, while the sparse AGP performed poorly.
-Table LABEL:tab:sm:fusion in the appendix presents the results for parameter estimation under the correct specification of the the HGP model.
+Table S.3 in the Appendix Material presents the results for parameter estimation under the correct specification of the the HGP model.
 We used the same metrics as in Section 4.1 to assess the parameters estimation.
@@
 Unlike the areal simulation, the coverage of some of the variance parameters was consistently below the nominal level of $95\%$.
-This results are expected, as the HGP inherits the weak identifiability of these parameters from geostatistical models under a normal likelihood (Zhang, 2004).
+This results are expected, as the HGP inherits the weak identifiability of these parameters from geostatistical models under a normal likelihood (Zhang,, 2004).
 Predictions remain trustworthy despite this issue.
@@
 We first analyze a dataset on the count of hospital admissions due to respiratory disease in 2010 for the 134 intermediate zones (IZ) to the north of the river Clyde in the Great Glasgow and Clyde health board in Scotland.
-This dataset is publicly available in theCARBayesdata package (Lee, 2022).
+This dataset is publicly available in theCARBayesdata package (Lee,, 2022).
 For each IZ, the expected number of admissions based on the national age- and sex-standardized rates, alongside the percentage of people classified as income deprived in each IZ are also available.
-The standardized morbidity ratio (SMR), the ratio between the number of observed and expected cases, is often used to control for confounding factors when estimating risks (Greenland, 1984).
+The standardized morbidity ratio (SMR), the ratio between the number of observed and expected cases, is often used to control for confounding factors when estimating risks (Greenland,, 1984).
 The goals of the analyses are: (1) to provide smooth estimates of the SMR, adjusted for the percentage of income deprivation, and (2) to contrast the outcomes from the proposed methodology with other established methods.
@@
 We analyze the data using the GLMM model from Section 3.1, with a Poisson likelihood, a $\log$ link function, an intercept, a covariate $x_{i}$, and an offset $E_{i}$.
-We employed the homoscedastic HGP (with a PEC function with smoothness $\nu=0.7$) as the prior for the random effects; the choice of $\nu=0.7$ is justified in Table LABEL:tab:sm:gof_scot of the appendix.
+We employed the homoscedastic HGP (with a PEC function with smoothness $\nu=0.7$) as the prior for the random effects; the choice of $\nu=0.7$ is justified in Table S.4 of the Appendix Material.
 The priors for the parameters in the HGP GLMM were set to be those presented in Section 3.2.
@@
 The highest split-$\hat{R}$ was smaller than $1.01$, providing evidence of convergence.
-Two competing models, BYM and DAGAR, were considered as the priors for the spatial random effects $\mathbf{z}$ for comparison, as outlined in Sections 4.1 and LABEL:sec:sm:sim, respectively.
-In the BYM model, parametrized by the marginal precision (Riebler et al., 2016), we followedMorris et al.
-(2019) and set standard normal priors for the intercept ($\beta_{0}$), regression coefficient ($\beta_{1}$), and the $\mathrm{logit}$ of the mixing parameter $\zeta$.
+Two competing models, BYM and DAGAR, were considered as the priors for the spatial random effects $\mathbf{z}$ for comparison, as outlined in Sections 4.1 and S.2.1, respectively.
+In the BYM model, parametrized by the marginal precision (Riebler et al.,, 2016), we followed Morris et al., (2019) and set standard normal priors for the intercept ($\beta_{0}$), regression coefficient ($\beta_{1}$), and the $\mathrm{logit}$ of the mixing parameter $\xi$.
 Similarly, a standard normal distribution truncated on positive real numbers was used for $\sigma$.
@@
 For these models, the parameters split-$\hat{R}$ provided evidence of convergence.
-In addition, trace and density plots for posterior samples of all fitted model parameters are provided in Section LABEL:sec:sm:areal_convergence of the appendix.
+In addition, trace and density plots for posterior samples of all fitted model parameters are provided in Section S.5.1 of the Appendix Material.
 Table 2: Parameter estimates and model comparison criteria for the three models fitted to the areal dataset. Marginal posterior summary through median and 95% credible intervals.
@@
 In other words, this credible interval suggests that the clustering effect ranges from almost negligible to very strong.
-A similar pattern is observed for the $\zeta$ in the BYM model.
+A similar pattern is observed for the parameter $\xi$ in the BYM model.
 Figure 5: Density histogram of the pairwise Hausdorff distances observed in the dataset and. In blue, estimated posterior density of the practical range $\rho$, and, in red, the prior used for $\rho$.
@@
 Notably, the majority of the posterior distribution mass for $\rho$ falls below the concentration of most pairwise distances, indicating a negligible clustering effect in hospital admissions.
-An additional analysis of the spatial correlation is presented in Section LABEL:sec:sm:areal_comp in the appendix .
+An additional analysis of the spatial correlation is presented in Section S.5 in the Appendix Material.
 In Table 2, we also present two model comparison criteria described in Section 4 for the models fitted to the data.
@@
 However, none of the models delivered a remarkably better fit to this dataset.
-This is further reinforced by looking at the maps of observed and smoothed SMR estimates presented in Figure LABEL:fig:sm:smr from the appendix.
+This is further reinforced by looking at the maps of observed and smoothed SMR estimates presented in Figure S.3 from the Appendix Material.
 5.2 Data fusion: Air pollution in Ventura and Los Angeles
@@
 Our second application analyzes fused fine particulate matter (PM2.5) data in Ventura and Los Angeles counties, California, USA.
-The dataset combines point-referenced direct measurements (EPA, 2018) from 1999 to present with areal satellite-derived estimates (Van Donkelaar et al., 2015), averaged from 2010 to 2012 (Figure 5).
+The dataset combines point-referenced direct measurements (U.S.
+Environmental Protection Agency Clean Air Markets Division,, 2018) from 1999 to present with areal satellite-derived estimates (Van Donkelaar et al.,, 2015), averaged from 2010 to 2012 (Figure 6).
 To align temporally, we averaged the direct measurements over the same period.
@@
 In this case, there are no covariates, the likelihood is Normal with variance $\tau^{2}$, and the link function is the identity.
-A heteroscedastic HGP prior is placed on the random effect allowing for differing variances between point-referenced and areal data through the SD function from Equation (2) with $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)\}$.
+A heteroscedastic HGP prior is placed on the random effect allowing for differing variances between point-referenced and areal data through the SD function from Equation (2) with $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$.
 For interpretability, we present results for $\sigma=\exp\{\alpha_{0}\}$ and $\sigma_{a}=\exp\{\alpha_{0}+\alpha_{1}\}$ instead of $\alpha_{0}$ and $\alpha_{1}$.
 The model was fit with $\nu$ fixed at different values, and we selected the model with best goodness-of-fit ($\nu=0.8$).
-For more details, see Table LABEL:tab:sm:gof_la in the appendix.
+For more details, see Table S.5 in the Appendix Material.
 The priors for model parameters are as in Section 4.2.
@@
 All model parameters exhibited split-$\hat{R}$ values below, indicating good convergence.
-This is further supported by the trace plots presented in Section LABEL:sec:sm:df_convergence of the appendix.
+This is further supported by the trace plots presented in Section S.6.1 of the Appendix Material.
 For comparison, we also fit an AGP random effect model, with priors for the model parameters as described in Section 4.2.
-Two meshes of different resolutions were considered; see Figure LABEL:fig:sm:meshes in the appendix.
+Two meshes of different resolutions were considered; see Figure S.17 in the Appendix Material.
 The sparse- and fine-mesh AGP models are denoted AGP1 and AGP2, respectively.
@@
 While the AGP2 achieves similar coverage, its prediction intervals are considerably wider, resulting in inferior interval scores.
-Further analysis by data type is detailed in Table LABEL:tab:sm:out-pred_df of the appendix, revealing the unreliability of competing methods for point-referenced data, where frequentist coverage can be as low as 3%.
+Table S.6 in the Appendix Material provides a breakdown by data type, highlighting the poor performance of competing methods for point-referenced data.
+The PI frequentist coverage, for example, can be as low as 3%.
 The drastic differences in the AGP results for different mesh resolutions are striking.
@@
 This weak dependence may imply AGP2 assigns nearly independent random intercepts to each data point, with variances inversely proportional to the sample unit’s area.
-AGP mesh resolution is typically practitioner-informed, and results are highly sensitive to this tuning parameter, which lacks a consensus on “optimal resolution” (Fuentes and Raftery, 2005; Liu et al., 2011; Wang and Furrer, 2021).
+AGP mesh resolution is typically practitioner-informed, and results are highly sensitive to this tuning parameter, which lacks a consensus on “optimal resolution” (Fuentes and Raftery,, 2005; Liu et al.,, 2011; Wang and Furrer,, 2021).
 Figure 7: Change-of-support of PM2.5 to a finer resolution using the sparse and fine AGP and the HGP priors.
@@
 The proposed framework unifies the modeling of spatial data, extending geostatistical models to include areal and fused data within the same setting.
-We established conditions under which the PEC function of the HGP is positive definite, ensuring its validity.
 As a prior for spatial random effects in a GLMM, the HGP proves to be competitive to specialized models for diverse spatial data types.
@@
 A crucial focus should be developing correlation functions designed specifically for the HGP, particularly those suited to scenarios where spatial data is observed on a sphere.
-Approximations inspired by methods like the Nearest Neighbor Gaussian Process (Datta et al., 2016; Quiroz et al., 2023) or reduced rank approximations (Solin and Särkkä, 2020) could significantly broaden the applicability of the HGP to large-scale datasets.
+Approximations inspired by methods like the Nearest Neighbor Gaussian Process (Datta et al.,, 2016; Quiroz et al.,, 2023) or reduced rank approximations (Solin and Särkkä,, 2020) could significantly broaden the applicability of the HGP to large-scale datasets.
 These advancements hold the potential to establish the HGP as an even more powerful and versatile tool for spatial data analysis.

### v4

@@
-Belo Horizonte, MG, Brazil
+Statistical Inferences and Predictions for Areal Data and Spatial Data Fusion with Hausdorff–Gaussian Processes
 Abstract
-Accurate modeling of spatial dependence is pivotal in analyzing spatial data, influencing parameter estimation and predictions.
-The spatial structure of the data significantly impacts valid statistical inference.
-Existing models for areal data often rely on adjacency matrices, struggling to differentiate between polygons of varying sizes and shapes.
-Conversely, data fusion models rely on computationally intensive numerical integrals, presenting challenges for moderately large datasets.
+Accurate modeling of spatial dependence is pivotal for both parameter estimation and prediction in spatial data analysis.
+The spatial structure of the data significantly narrows the set of available models.
+Existing areal data models often ignore polygon geometry, while data fusion models require computationally intensive integration, limiting their application.
 In response to these issues, we propose the Hausdorff-Gaussian process (HGP), a versatile model utilizing the Hausdorff distance to capture spatial dependence in both point and areal data.
@@
 We validate our approach through a comprehensive simulation study and application to two real-world scenarios: one involving areal data and another demonstrating its effectiveness in data fusion.
-The results suggest that the HGP is competitive with specialized models regarding goodness-of-fit and prediction performances.
-In summary, the HGP offers a flexible and robust solution for modeling spatial data of various types and shapes, with potential applications spanning fields such as public health and climate science.
-Keywords: CAR models; Geostatistics; Spatial data fusion; Spatial modeling; Spatial prediction
+Results show the HGP is a competitive, flexible, and robust solution for modeling diverse spatial data, with potential applications in fields like public health and climate science.
+Keywords: CAR models, Geostatistics, Spatial data fusion, Spatial modeling, Spatial prediction
 1 Introduction
 Accounting for spatial dependence is pivotal in spatial data analysis.
-Valid statistical inference are deeply influenced by the spatial structure and geometry of the data.
-Point-referenced data are typically analyzed using geostatistics methods (e.g., Cressie,, 1993).
-For areal data, the conditional autoregressive model (Besag,, 1974, CAR) and its variants (Besag et al.,, 1991; Rodrigues and Assunçao,, 2012; Datta et al.,, 2019; Cruz-Reyes et al.,, 2023) are often preferred.
 The nature of the observed spatial data significantly influences the choice of statistical methodology.
-This becomes increasingly critical in complex inferential problems, such as predicting a variable in a target map using data from a source map.
-This issue, known as change of support (Cressie,, 1993, pg.
-284–289), requires tailored methods depending on the spatial units involved (Gelfand et al.,, 2001).
-Spatial data from diverse sources often differ in spatial units or supports, resulting in spatial misalignment (Gotway and Young,, 2002).
-Data fusion is necessitated when a spatial variable is provided by multiple, potentially misaligned, sources (Moraga et al.,, 2017).
+Point-referenced data are typically analyzed using geostatistics methods (e.g., cressie1993statistics).
+For areal data, the conditional autoregressive model (besag1974spatial, CAR) and its variants (besag1991bayesian; rodrigues2012bayesian; datta2019spatial; cruz2023inducing) are often preferred.
+These differences become increasingly critical in complex inferential problems, such as predicting a variable in a target map using data from a source map.
+This issue, known as change of support (cressie1993statistics, pg.
+284–289), requires tailored methods depending on the spatial units involved (gelfand2001change).
+Spatial data from diverse sources often differ in spatial units or supports, resulting in spatial misalignment (gotway2002combining).
+Data fusion is necessitated when a spatial variable is provided by multiple, potentially misaligned, sources (moraga2017geostatistical).
 Both traditional areal models and data fusion models for spatial data exhibit inherent limitations.
-Traditional areal models face challenges with predictions for overlapping spatial units (Bakar and Jin,, 2020), do not account for region size or shape of the areal units and lack interpretable spatial dependence parameters (Assunção and Krainski,, 2009; Cruz-Reyes et al.,, 2023).
-An exception of the last limitation is the directed acyclic graph auto-regressive (DAGAR) model, which offers an interpretable parameter (Datta et al.,, 2019).
-Data fusion models, which view areal data as averages of a continuous process, may be unsuitable for discrete outcomes (Gelfand et al.,, 2001).
-Additionally, their reliance on arbitrary discretization for likelihood evaluation lacks consensus on optimal resolution (Fuentes and Raftery,, 2005; Liu et al.,, 2011; Wang and Furrer,, 2021) and can introduce unquantifiable biases (Gonçalves and Gamerman,, 2018).
-We introduce a unified and flexible approach for analyzing spatial data of varying types, encompassing areal, point-referenced, and mixed spatial supports.
-Our key objective is to develop a model capable of seamlessly handling these diverse data formats within a single framework.
-This is achieved by combining Gaussian processes (GPs) with the Hausdorff distance (Arutyunov et al.,, 2016).
+Traditional areal models lack interpretable spatial dependence parameters (assuncao2009neighborhood).
+An exception of that last limitation is the directed acyclic graph auto-regressive (DAGAR) model, which offers an interpretable parameter (datta2019spatial).
+These models, which rely on a fixed adjacency graph, are not suited for out-of-sample prediction, as introducing new spatial units fundamentally alters the whole model structure.
+Furthermore, when overlapping spatial units are present (bakar2020areal), another limitation is that they do not account for the size or shape of the areal units.
+Data fusion models, which view areal data as averages of a continuous process, may be unsuitable for discrete outcomes (gelfand2001change).
+Additionally, their reliance on arbitrary discretization for likelihood evaluation lacks consensus on optimal resolution (fuentes2005model; liu2011empirical; wang2021combining) and can introduce unquantifiable biases (goncalves2018exact).
+We introduce a unified approach for areal, point-referenced, and mixed spatial data by combining Gaussian processes (GPs) with the Hausdorff distance (arutyunov2016some), which seamlessly handles these diverse formats in a single framework.
 The Hausdorff distance, defined on a reference metric space, effectively measures the distance between two sets while accounting for their shape.
 For point pairs, it reduces to the distance metric associated to the reference metric space.
-GPs, recognized for their flexibility, are widely utilized in spatial data modeling, notably as priors in generalized mixed effects models (Diggle et al.,, 1998), in spatially varying regression coefficients (Gelfand et al.,, 2003), and in addressing spatial misalignment issues (Banerjee and Gelfand,, 2002).
+GPs, recognized for their flexibility, are widely utilized in spatial data modeling, notably as priors in generalized mixed effects models (diggle1998model), in spatially varying regression coefficients (gelfand2003spatial), generalized multilevel models (prates2024fast), and in addressing spatial misalignment issues (banerjee2002prediction).
 We propose the Hausdorff-Gaussian process (HGP), a GP with a correlation function defined on the Hausdorff distance between spatial units.
@@
 A crucial aspect of our work is to assess the competitiveness of the HGP, particularly when compared to specialized models designed for areal and mixed spatial data.
-We investigate whether the HGP, using a powered exponential correlation, can effectively model areal and mixed spatial data, thereby offering a novel and versatile tool for analyzing spatial and spatiotemporal data with mixed supports.
+We investigate whether the HGP, using a powered exponential correlation, can effectively model areal and mixed spatial data, thereby offering a novel and versatile tool for analyzing spatial data with mixed supports.
 This paper is structured as follows.
 Section 2 introduces the Hausdorff distance, discussing its properties, and then delineates the HGP, highlighting its significant applications.
-Section 3 sets up spatial generalized linear mixed models (GLMM) with HGP components and details their inferences in the Bayesian framework.
+Then, in Section 3, we set up spatial generalized linear mixed models (GLMM) with HGP components and detail their inferences in the Bayesian framework.
 A simulation study to compare the proposed model with specific models for areal and fused data is presented in Section 4.
@@
 Section 6 concludes with a discussion.
+For completeness, we introduce models designed specifically for areal and fused data in the Appendix.
+The Supplementary Material and the code to reproduce the analyses are available in the following GitHub repository: https://github.com/lcgodoy/jabe-d-25-00205.
 2 Hausdorff–Gaussian Process
@@
 Specifically, it is the greatest distance from a point in one set to the closest point in the other set.
-Consider a metric space $(D,d)$, where $D$ is a spatial region of interest and $d(x,y)$ defines the distance between any two elements $x,y\in D$.
+Consider a metric space $(D,d)$, where $D$ is a spatial region of interest and $d(x,y)$ defines a distance between any two elements $x,y\in D$.
 For a single point $x$ and a set $A$, define $d(x,A)=\inf_{a\in A}d(x,a)$ as the closest distance from $x$ to any element in $A$.
-The directed Hausdorff distance from a set $A$ to a set $B$ is
-$\vec{h}(A,B)=\sup_{a\in A}d(a,B).$
-This measure is not necessarily symmetric. The symmetric Hausdorff distance is defined as the greater of the two directed Hausdorff distances:
+The directed Hausdorff distance from a set $A$ to a set $B$ is $\vec{h}(A,B)=\sup_{a\in A}d(a,B)$.
+This measure is not necessarily symmetric.
+The symmetric Hausdorff distance is defined as the greater of the two directed Hausdorff distances:
 $h(A,B)=\max\left\{\vec{h}(A,B),\vec{h}(B,A)\right\}.$
-Note that if $A$ and $B$ are both singletons, then $h(A,B)=d(A,B)$. Interestingly, if $A$ and $B$ are compact, there exist $a\in A$ and $b\in B$ such that $h(A,B)=d(a,b)$ (Arutyunov et al.,, 2016).
-In spatial analysis, the Hausdorff distance provides a robust and suitable metric over non-empty compact subsets of $D$, denoted as $\mathcal{B}(D)$ (Molchanov,, 2005, Appendix C).
-While not functioning as a conventional distance in unbounded sets, it is highly effective in bounded contexts, which are prevalent in spatial statistics.
-Typically, the study region $D$ in spatial statistics is a bounded subset of $\mathbb{R}^{2}$, where the Hausdorff distance excels in quantifying distances between spatial units by capturing both their shapes and sizes.
+Note that if $A$ and $B$ are both singletons, then $h(A,B)=d(A,B)$. If they are compact, there exist $a\in A$ and $b\in B$ such that $h(A,B)=d(a,b)$ (arutyunov2016some).
+In spatial statistics, sample units (e.g., points, polygons) are repsented as non-empty, closed, and bounded subsets of a study region $D\subset\mathbb{R}^{k}$.
+In Euclidean domains, the Heine-Borel theorem asserts that closed and bounded sets are compact (abbott2015understanding, pg.
+98).
+The Hausdorff distance is a suitable metric for the collection of these non-empty compact subsets, hereafter denoted $\mathcal{B}(D)$ (molchanov2005theory, Appendix C).
+By the properties of a metric, $h(A,B)=0$ if and only if $A$ and $B$ are identical.
 The metric space induced by the tuple $(\mathcal{B}(D),h)$ inherits properties like compactness and completeness from $(D,d)$.
-Further details regarding the topology induced by the Hausdorff distance can be found in Appendix C inMolchanov, (2005).
-AB
-AB
-A.B
-Figure 1: Three different pairs of bounded subsets of the Euclidean plane.
-In (a) the sets AA and $B$ share a single point on their border.
-In (b), they have non-empty intersection.
-While in (c), $B$ is a singleton set contained by $A$.
-The dashed lines represent the Hausdorff distance between the sets.
-To demonstrate the efficacy of the Hausdorff distance in measuring distances between various spatial units, Figure 1 presents three simple examples.
-In each sub-figure, set $A$ is represented as a circle with a radius of 1.
-Figure 1a introduces a set $B$, a circle tangent to $A$.
-Figure 1b examines the distance between two sets having a non-empty intersection.
-Figure 1c features $B$ as an interior point of $A$.
-The dashed lines in the figures delineate the Hausdorff distances between the sets, highlighting the metric’s ability to consider the shape, size, and orientation of the objects.
-This contrasts with other simpler distance functions between non-empty sets, such as the infimum distance between two sets, denoted $\inf_{x\in X}\inf_{y\in Y}d(x,y)$.
-This distance does not fulfill the metric property of positiveness condition.
-In other words, $d(x,y)=0$ does not imply $x=y$.
-Indeed, employing such distance latter would yield a distance of zero between the three pairs of sets depicted in Figure 1.
+Further details regarding the topology induced by the Hausdorff distance can be found in Appendix C in molchanov2005theory.
+Lastly, in the Supplementary Material, Section S.1, we provide a toy example comparing the Hausdorff distance to the distance between centroids and, in addition, between borders.
 2.2 Hausdorff–Gaussian Process
@@
 A GP is a stochastic process over an index set such that all its finite-dimensional marginal distributions are multivariate Normal distributions.
-An HGP expands the index set on a spatial region $D$ to encompasses a broader range of spatial units, $\mathcal{B}(D)$, which contains non-empty, bounded, and compact subsets of $D$.
-To define an isotropic HGP $Z(\cdot)$ over $\mathcal{B}(D)$, we specify a marginal mean function function $m(\cdot)$, a positive-valued marginal standard deviation (SD) function $v(\cdot)$, and an isotropic correlation function $r(h)$ for any two sets apart in Hausdorff distance $h$.
-With these functions, an HGP over $\mathcal{B}(D)$ is a stochastic process $Z(\cdot)$ whose marginal distribution at $\mathbf{s}_{1},\ldots,\mathbf{s}_{p}\in\mathcal{B}(D)$ is multivariate normal with mean vector and covariance matrix, respectively,
+Let $D$ be a spatial region of interest.
+An HGP index set is taken to be the class of non-empty compact subsets of $D$, denoted $\mathcal{B}(D)$.
+This class encompasses both singletons and polygons, which can be represented as closed and bounded sets.
+To define an isotropic HGP $Z(\cdot)$ over $\mathcal{B}(D)$, we specify a marginal mean function $m(\cdot)$, a positive-valued marginal standard deviation (SD) function $v(\cdot)$, and a valid isotropic correlation function $r(h)$ for any two sets apart in Hausdorff distance $h$.
+With these functions, an HGP is a stochastic process $Z(\cdot)$ whose marginal distribution at $\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}\in\mathcal{B}(D)$ is multivariate Normal with mean vector and covariance matrix, respectively,
 $\mathbbm{E}[Z(\mathbf{s}_{i})]=m(\mathbf{s}_{i})\quad\text{and}\quad\mathrm{Cov}[Z(\mathbf{s}_{i}),Z(\mathbf{s}_{j})]=v(\mathbf{s}_{i})v(\mathbf{s}_{j})r\big(h(\mathbf{s}_{i},\mathbf{s}_{j})\big),$ (1)
-for $i,j=1,\ldots,p$, where $h(\mathbf{s}_{i},\mathbf{s}_{j})$ is the Hausdorff distance between $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$.
+for $i,j=1,\ldots,n$, where $h(\mathbf{s}_{i},\mathbf{s}_{j})$ is the Hausdorff distance between $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$.
 This process allows for representations of diverse spatial data, from areal to mixed, making it suitable for a wide range of applications such as disease mapping and data fusion.
 Denote the HGP process defined in Equation (1) by $\textrm{HGP}\{m(\cdot),v(\cdot),r(\cdot)\}$.
-The mean function $m(\cdot)$ captures large-scale variations, usually through covariates.
-The SD function $v(\cdot)$ function accommodates both homoscedastic and hereroscedastic scenarios.
-For instance, with a log link and covariate $w(\mathbf{s})$ at spatial unit $\mathbf{s}$, a natural model is
-$\log v(\mathbf{s})=\alpha_{0}+\alpha_{1}w(\mathbf{s}).$ (2)
-Equation (2) allows a distinction between point-referenced and areal data, a need in data fusion applications (Wang and Furrer,, 2021).
-With $\mathcal{A}(\cdot)$ being the function that returns the area of a spatial object.
-A covariate $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$ enables different variances between these data types.
-Certainly, care is necessary in specifying $v(\mathbf{s})$ to ensure a valid GP (Palacios and Steel,, 2006).
+A GP’s mean function has little impact on the process’ local behavior (stein1999interpolation, pg.
+12).
+They capture large-scale variations, and, hereafter $m(\cdot)$ is either held constant or defined through covariates.
+The SD function $v(\cdot)$ accommodates both homoscedastic and hereroscedastic scenarios.
+For instance, with a log link, a natural model is
+$\log v(\mathbf{s})=\alpha_{0}+\alpha_{1}w(\mathbf{s}),$ (2)
+where $w(\mathbf{s})$ is a covariate available for any $s\in\mathcal{B}(D)$.
+Equation (2) allows a distinction between point-referenced and areal data, a need in data fusion applications (wang2021combining).
+Specifically, to enable different variances between these data types, we can define $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$, where $\mathcal{A}(\cdot)$ is a function that returns the area of a spatial object.
+Certainly, care is necessary in specifying $v(\mathbf{s})$ to ensure a valid GP (palacios2006non).
 2.3 Correlation Function
-The correlation function $r(\cdot)$ needs to be carefully chosen to ensure the validity of the HGP as it defines the spatial dependence structure across different locations.
+The correlation function $r(\cdot)$ needs to be carefully chosen to ensure the validity of the HGP.
 In particular, we focus on isotropic correlation functions.
-Therefore, unless explicitly mentioned, all subsequent references to correlation functions will assume they are isotropic.
-A valid correlation function must be positive definite and meet two key criteria: $r(0)=1$ and $\lim_{h\to\infty}r(h)=0$ (Cressie,, 1993, Ch.
+Thus, unless explicitly mentioned, all subsequent references to correlation functions will assume they are isotropic.
+A valid correlation function must be positive definite and meet two key criteria: $r(0)=1$ and $\lim_{h\to\infty}r(h)=0$ (cressie1993statistics, Ch.
 2.3.2), where $h$ is the Hausdorff distance between a pair $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{B}(D)$.
 The latter ensures that spatial correlation diminishes with increasing distance.
-It is noteworthy that while most geostatistical correlation functions are constructed for Euclidean spaces (Cressie,, 1993, Ch.
-2.5.1), their positive definiteness may not extend to non-Euclidean spaces (Curriero,, 2006; Li et al.,, 2023).
+It is noteworthy that while most geostatistical correlation functions are constructed for Euclidean spaces (cressie1993statistics, Ch.
+2.5.1), their positive definiteness may not extend to non-Euclidean spaces (curriero2006use).
 For practical purposes, we propose to use a two-parameter powered exponential correlation (PEC) function
@@
 Consequently, $\rho$ is the distance at which the spatial correlation between a pair of spatial units drops to $0.10$, commonly referred to as the practical range.
-The PEC function is appealing because, for $\nu\in(0,1]$, it is positive definite as long as the distance function is conditionally negative definite (Schoenberg,, 1938).
-To the best of our knowledge, there exist no theoretical results on positive definite functions for the Hausdorff distance.
-Therefore, we recommend an empirical assessment of the PEC function positive definiteness before analyzing the data.
+The PEC function is appealing because, for $\nu\in(0,1]$, it is positive definite as long as applied to a conditionally negative definite (CND) distance function (schoenberg1938metric).
+However, to the best of our knowledge, no theoretical results establish when the Hausdorff distance is CND, nor are there results concerning positive definite functions of this distance.
+Thus, we recommend empirically assessing the positive definiteness of the PEC function before analyzing the data.
 To this end, we empirically assess the positive definiteness of the PEC function for $\nu\in(0,1]$.
 We first calculate Hausdorff distance matrices for the areal and fused data to be analyzed in Sections 5.1 and 5.2, respectively.
-Evaluating the PEC function across these matrices with $\nu\in[0.01,1]$ and $\rho\in[d_{\min},d_{\max}]$ (where $d_{\min}$ and $d_{\max}$ are the dataset’s minimum and maximum distances) yields corresponding correlation matrices.
-We then compute each matrix’s smallest eigenvalue.
-A negative eigenvalue suggests the corresponding PEC function is not positive definite.
-However, even valid covariance functions can be ill-conditioned or nearly rank-deficient at dense locations or when there exists a severely strong spatial dependence (Banerjee et al.,, 2012).
+For comparison purposes, we re-scale these distance matrices so that all their elements are less than or equal to 1.
+We then evaluate the PEC function for these matrices using a grid of values for $\nu\in(0,1]$ and $\rho\in(0,1.5]$, where the upper bound of $\rho$ represents 1.5 times the largest distance observed in the respective matrix.
+This process generates a set of corresponding correlation matrices.
+Finally, we compute the smallest eigenvalue for each of these correlation matrices.
+A negative eigenvalue indicates the corresponding PEC function is not positive definite for that combination of parameters.
+However, even valid covariance functions can be ill-conditioned or nearly rank-deficient at dense locations or when there exists a strong spatial dependence (banerjee2013efficient).
 Figure 1: Smallest eigenvalue of the powered exponential correlation matrix under different combinations of smoothness ($\nu$) and practical range ($\rho$) for areal and fused data.
 The transparency indicates regions where $\lambda_{1}<0$.
-Figure 2 displays the smallest eigenvalues ($\lambda_{1}$) of the PEC-induced correlation matrices for both datasets.
+Figure 1 displays the smallest eigenvalues ($\lambda_{1}$) of the PEC-induced correlation matrices for the fused and areal datasets from Sections 5.1 and 5.2, respectively.
 The fused data exhibits positive eigenvalues for all $(\nu,\rho)$ pairs.
-Conversely, the areal data reveals some small negative eigenvalues (range 0.0099-0.0099 to 0.00006-0.00006) under severely strong spatial dependence and at $\nu\approx 1$, the border of the parametric space.
-Therefore, our empirical study suggests the positive definiteness of the PEC function for $\nu\in(0,1]$.
+In contrast, the areal data reveals some small negative eigenvalues (range 0.0099-0.0099 to 0.00006-0.00006) under strong spatial dependence and at $\nu\approx 1$, the border of the parametric space.
+Based on these empirical results, it is safe to use the PEC function with $\nu\in(0,1]$ for fusion data, but, for the areal application, it may be advisable to stick with a $\nu$ value less than 1.
+We further investigate the PD property of the PEC function across different datasets and domains.
+Similiar conclusion are obtained and these results are presented in Section S.2 of the Supplementary Material.
 3 HGP in Spatial GLMM
@@
 Generalized linear models for spatial data can benefit from replacing the GP with the HGP as spatial random effects.
-In this context, assume $\mathcal{S}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}$ is a set of spatial units or locations such that $\mathbf{s}_{i}\in\mathcal{B}(D)$, where $D\subset\rm I\!R^{2}$.
+In this context, assume $\mathcal{S}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}$ is a set of spatial units such that $\mathbf{s}_{i}\in\mathcal{B}(D)$, where $D\subset\rm I\!R^{2}$.
 Consider a response variable $Y(\mathbf{s}_{i})$ and a $p$-dimensional vector of covariates $\mathbf{X}_{i}$ associated to the $i$-th spatial unit.
 Let $\mathbf{Z}={\big[Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n})\big]}^{\top}$ be an $n$-dimensional vector of latent spatial random effects.
-Conditional on $Z(\mathbf{s}_{i})$ and $\mathbf{X}_{i}$, we assume that the responses $Y(\mathbf{s}_{i})$’s are mutually independent.
-The conditinal distribution of $Y(\mathbf{s}_{i})$ is characterized by its mean mean and possibly some additional parameters.
-This leads to a spatial GLMM built on an HGP.
-Specifically, the spatial GLMM is specified as
-$Y(\mathbf{s}_{i})\mid\mathbf{X}_{i},Z(\mathbf{s}_{i})\overset{{\rm ind}}{\sim}f(\cdot\mid\mu_{i},\bm{\gamma}),$ (3)
-where $f(\cdot\mid\mu_{i},\bm{\gamma})$ is a probability density or mass function of a distribution with expectation $\mu_{i}$ and an additional parameter vector $\bm{\gamma}$.
-The conditional mean $\mu_{i}$ is linked to the covariate vector and the spatial random effects through a suitable, known link function $g$:
+Then, the spatial GLMM is specified as
+$Y(\mathbf{s}_{i})\mid\mathbf{X}_{i},Z(\mathbf{s}_{i})\overset{{\rm ind}}{\sim}f(\cdot\mid\mu_{i},\bm{\gamma}),$
 $\mu_{i}=\mathbbm{E}[Y(\mathbf{s}_{i})\mid\mathbf{X}_{i},Z(\mathbf{s}_{i})]=g^{-1}(\mathbf{X}^{\top}_{i}\bm{\beta}+Z(\mathbf{s}_{i})),$ (3)
-where $\bm{\beta}$ is a $p$-dimensional regression coefficient vector.
-The spatial dependence among the responses $\mathbf{Y}={\big[Y(\mathbf{s}_{1}),\ldots,Y(\mathbf{s}_{n})\big]}^{\top}$ is introduced by an HGP on $\mathbf{Z}$, namely,
-$\mathbf{Z}\sim\textrm{HGP}\big\{0,v(\cdot;\bm{\sigma}),r(\cdot;\bm{\delta})\big\},$ (5)
-with zero-mean, a SD function $v(\cdot;\bm{\sigma})$ with parameters $\bm{\sigma}$, and the PEC function $r(\cdot,\bm{\delta})$ with parameters $\bm{\delta}=\{\rho,\nu\}$.
+where $g$ is a link function, $\bm{\beta}$ is a $p$-dimensional regression coefficient vector, and $f(\cdot\mid\mu_{i},\bm{\gamma})$ is a probability density or mass function of a distribution with expectation $\mu_{i}$ and an additional parameter vector $\bm{\gamma}$.
+Thus, conditional on $Z(\mathbf{s}_{i})$ and $\mathbf{X}_{i}$, we assume that the responses $Y(\mathbf{s}_{i})$’s are mutually independent.
+The spatial dependence among the responses $\mathbf{Y}={\big[Y(\mathbf{s}_{1}),\ldots,Y(\mathbf{s}_{n})\big]}^{\top}$ is introduced by a zero-mean HGP $\mathbf{Z}\sim\textrm{HGP}\big\{0,v(\cdot;\bm{\sigma}),r(\cdot;\bm{\delta})\big\}$, where $\bm{\sigma}$ and $\bm{\delta}=\{\rho,\nu\}$, are the SD and PEC functions’ parameters, respectively.
 To summarize, the parameters of the spatial GLMM are $\bm{\theta}={\{\bm{\beta}^{\top},\bm{\sigma}^{\top},\bm{\delta}^{\top},\bm{\gamma}^{\top}\}}^{\top}$.
-A general spatial GLMM is obtained by choosing a suitable $f(\cdot\mid\mu_{i},\bm{\gamma})$ for the response.
-For binary response variables, a Bernoulli distribution yields a spatial logistic regression.
-Count data can be modeled with either a Poisson or Negative Binomial distribution, depending on the presence of overdispersion.
-When the conditional distribution of the response is normal, the additional parameter $\gamma$ is the measurement error or small scale variance parameter.
-Let $\mathbf{X}$ be a $n\times p$ matrix with its $i$-th row given by $\mathbf{X}_{i}$.
-From the conditional dependence, the joint density (or mass) function of $\mathbf{Y}$ conditioning on $\mathbf{X}$ and $\mathbf{Z}$ is
-$p(\mathbf{y}\mid\mathbf{X},\mathbf{z})=\prod_{i=1}^{n}f(y(\mathbf{s}_{i})\mid\mu_{i},\bm{\gamma}),$ (6)
-where $\mathbf{y}$ and $\mathbf{z}$ are realizations of $\mathbf{Y}$ and $\mathbf{Z}$, respectively.
 3.2 Bayesian Inferences
 Prior distributions for all the model parameters are to be specified to complete the Bayesian model specification.
-For fixed effects regression coefficients ($\bm{\beta}$), we employ uncorrelated zero-mean Normal priors with marginal variances $10^{2}$.
-In the homoscedastic case, on $\sigma$ we place a half-t prior with $3$ degrees of freedom, denoted $t_{+}(3)$ (Gelman,, 2006).
-For the hereroscedastic model (2), we set independent standard normal priors for $\alpha_{0}$ and $\alpha_{1}$.
+For fixed effects regression coefficients ($\bm{\beta}$), we employ uncorrelated zero-mean Normal priors with marginal variances $\mathrm{v}_{\beta}$.
+In the homoscedastic case, on $\sigma$ we place a half-t prior with $3$ degrees of freedom, denoted $t_{+}(3)$ (gelman2006prior).
+For the hereroscedastic model (2), we set independent Normal priors for $\alpha_{0}$ and $\alpha_{1}$ with marginal variances $\mathrm{v}_{\alpha}$.
+Hereafter, we set the marginal variances as $\mathrm{v}_{\beta}=10^{2}$ and $\mathrm{v}_{\alpha}=1$ unless otherwise stated.
 The PEC dependence is encoded through $\bm{\delta}=\{\rho,\nu\}$.
-In geostatistics, smoothness parameters for correlation functions are typically fixed, and their estimation properties are rarely reported (Zhang,, 2004).
-Due to this and the scope of this paper, we specify separate models for different values of $\nu$ ensuring a positive definite PEC and select the best fit for inference.
-The spatial dependence parameter $\rho$ is also inherently difficult.
-Thus, to avoid oversmoothing, we employ an exponential prior for $\rho$ with mean $\lambda_{\rho}=-\rho_{0}/\log(p_{\rho})$, where $\rho_{0}$ is an upper-bound set to four-fifths of the largest observed Hausdorff distance, and $p_{\rho}$ is the probability that $\rho$ exceeds $\rho_{0}$.
+In geostatistics, smoothness parameters for correlation functions are typically fixed (zhang2004inconsistent; datta2016hierarchical).
+Due to this and the scope of this paper, we specify separate models for different values of $\nu$, selected based on the empirical assessment to ensure positive definiteness of the induced correlation matrices, and select the best fit for inference.
+The spatial dependence parameter $\rho$ is also inherently difficult to estimate.
+Therefore, to avoid oversmoothing, we set an exponential prior in the same spirit of the penalized complexity (PC) prior (simpson2017penalising; fuglstad2019constructing) for $\rho$.
+Specifically, we employ an exponential prior for $\rho$ with mean $\lambda_{\rho}=-\rho_{0}/\log(p_{\rho})$, where $\rho_{0}$ is an upper-bound set to four-fifths of the largest observed Hausdorff distance, and $p_{\rho}$ is the probability that $\rho$ exceeds $\rho_{0}$.
 The resulting prior concentrates its mass on lower values $\rho$.
+Note that, the resulting prior is not exactly a PC prior since it is not placed on the Kullback-Leibler divergence with a simpler model.
 The posterior distribution of $\bm{\theta}$ and $\mathbf{Z}$ given the observed data $\mathbf{y}$ and $\mathbf{X}$ is
-$\pi(\bm{\theta},\mathbf{z}\mid\mathbf{y})\propto p(\mathbf{y}\mid\mathbf{X},\mathbf{z},\bm{\beta},\bm{\gamma})\pi(\mathbf{z}\mid\bm{\sigma},\bm{\delta})\pi(\bm{\theta}),$ (7)
+$\pi(\bm{\theta},\mathbf{z}\mid\mathbf{y})\propto\pi(\mathbf{y}\mid\mathbf{X},\mathbf{z},\bm{\beta},\bm{\gamma})\pi(\mathbf{z}\mid\bm{\sigma},\bm{\delta})\pi(\bm{\theta}),$ (4)
 where $\pi(\cdot)$ denotes prior and posterior distributions of its arguments.
 Importantly, the distribution placed on the random effects $\mathbf{z}$ also functions as a prior, as we draw samples from these latent effects distribution conditioned on the observed data.
-Parameter Estimation
-The estimation of the spatial GLMM parameters is obtained from the posterior distribution defined in Equation (7).
-As the posterior distribution is intractable, we use the No-U-Turn (Homan and Gelman,, 2014) Markov Chain Monte Carlo (MCMC) sampler available in Stan (Team,, 2023) to draw samples from it.
-The No-U-Turn algorithm samples all the model parameters jointly and efficiently exploits the parameter space using automatic differentiation.
-Additionally, it eliminates the need for hand-tuning, making it a highly convenient sampler.
-We initialize the parameters with samples from their respective prior distributions.
+The estimation of the spatial GLMM parameters are obtained from the posterior distribution defined in Equation (4).
+Since this distribution is intractable, we draw samples from it using the No-U-Turn Markov chain Monte Carlo (MCMC) sampler in Stan (hoffman2014no; stan2023rstan).
+Unless otherwise stated, we initialize the parameters with samples from their respective prior distributions.
 The latent random effects are initialized from a standard Gaussian distribution.
 The number of samples and the warm-up period of the MCMC algorithm are application dependent.
-We assess the convergence of the chains using the split-$\hat{R}$ diagnostic (Vehtari et al.,, 2021).
+We assess the convergence of the chains using the split-$\hat{R}$ diagnostic (vehtari2021rank).
 Finally, the parameters’ point estimates and 95% credible intervals (CI) are obtained as, respectively, the median and percentiles ($2.5$ and $97.5$, unless otherwise stated) of the marginal MCMC samples.
-Model Comparison
-The assessment of goodness-of-fit (GoF) in the simulation study and applications involved two model comparison criteria: the leave-one-out information criterion (Vehtari et al.,, 2017, LOOIC), and the widely applicable information criterion (Watanabe and Opper,, 2010, WAIC).
-For both of them, lower values indicate a better fit.
-We did not use the deviance information criterion due to its assumption of a small number of parameters relative to the sample size (McElreath,, 2016, Ch.
-7.4), which would require computationally intensive numerical integration of random effects in some of our models.
+We assess goodness-of-fit (GoF) using the leave-one-out (vehtari2017practical, LOOIC) and widely applicable (watanabe2010asymptotic, WAIC) information criteria, where lower values indicate a better fit.
 When comparing a single candidate model to the HGP, we utilize the difference in the GoF metric $G$ between the HGP and the given model, denoted by $\Delta_{G}$.
@@
 The competing methods depend on the type of application we are working with and will be mentioned in the upcoming sections.
-Predictions
-Predicting spatial phenomena accurately at unobserved locations is crucial in many applications, especially in change of support settings (Gelfand et al.,, 2001).
+Predicting spatial phenomena accurately at unobserved locations is crucial in many applications.
 The HGP prior for spatial random effects in GLMMs offers a powerful solution.
@@
 Assuming covariates from Equation (3) are available, we can generate predictions accounting for uncertainty using the posterior predictive distribution of $\mathbf{Y}^{\ast}$:
-$p(\mathbf{y}^{\ast}\mid\mathbf{y})=\int p(\mathbf{y}^{\ast}\mid\mathbf{z}^{\ast},\bm{\theta})p(\mathbf{z}^{\ast}\mid\mathbf{z},\bm{\theta})\pi(\bm{\theta}\mid\mathbf{y}){\rm d}\bm{\theta},$ (8)
+$\pi(\mathbf{y}^{\ast}\mid\mathbf{y})=\int\pi(\mathbf{y}^{\ast}\mid\mathbf{z}^{\ast},\bm{\theta})\pi(\mathbf{z}^{\ast}\mid\mathbf{z},\bm{\theta})\pi(\bm{\theta}\mid\mathbf{y})\,{\rm d}\mathbf{z}^{\ast}\,{\rm d}\mathbf{z}\,{\rm d}\bm{\theta},$ (5)
 where $\mathbf{z}^{\ast}={(z(\mathbf{s}^{\ast}_{1}),\ldots,z(\mathbf{s}^{\ast}_{m}))}^{\top}$ is a realization of the spatial random effect at the new locations, and $\bm{\theta}$ represents the model parameters.
-Details on sampling from this distribution are provided in Section S.1 of the Appendix Material.
+Details on sampling from this distribution are provided in Section S.3 of the Supplementary Material.
 4 Simulation Studies
 4.1 Areal Data
-We conducted a simulation study to assess the performance of the HGP relative to the DAGAR model, a state-of-the-art model designed explicitly for areal data (Datta et al.,, 2019), in modeling data generated from other models.
+We conducted a simulation study to assess the performance of the HGP relative to the DAGAR model, a state-of-the-art model designed explicitly for areal data (datta2019spatial), in modeling data generated from other models.
+For details about the DAGAR model, we refer to the Appendix.
 We grounded our simulations in a real-world context by utilizing the 134 Intermediate Zones (IZ) north of the River Clyde within the Greater Glasgow in Scotland (Figure S.5), mirroring the application detailed in the upcoming Section 5.1.
@@
 The intercept and regression coefficients associated with this explanatory variable were fixed at $4.28$ and $0.33$, respectively, to mirror the parameters estimated in the real-world application (Section 5.1).
-Two random effect structures were considered: the aggregated Gaussian process (AGP) and the Besag–York–Mollié (Besag et al.,, 1991, BYM), with model parameters corresponding to their fits to the real data.
+Two random effect structures were considered: the Aggregated Gaussian Process (AGP) and the Besag–York–Mollié (besag1991bayesian, BYM), with model parameters corresponding to their fits to the real data.
+Details about those models and their parameters are given in the Appendix, Section A.
 The AGP assumes a point-level GP aggregated at the areal unit level and is closer in structure to the HGP model.
@@
 The BYM model allows for spatially correlated random effects with a structure closer to that of the DAGAR model.
-We set its marginal standard deviation (Riebler et al.,, 2016), denoted $\sigma$, to 0.24 and the spatial dependence parameter ($\xi$) to 0.43, aligning with estimates from the real data analyses in Section 5.1.
-Section S.2.1 of the Appendix Material provides further details on the models used to generate data.
+We set its marginal standard deviation (riebler2016intuitive), denoted $\sigma$, to $0.24$ and the spatial dependence parameter ($\xi$) to $0.43$, aligning with estimates from the real data analyses in Section 5.1.
+In Section S.5.2 from the Supplementary Material, we detail how to simulate data from the AGP.
 We generated 200 datasets from the Poisson GLMM under each of the two distinct random effect structures.
-For each simulated dataset, we fit two Bayesian Poisson GLMMs, the first one using an homoscedastic HGP while the second one using a DAGAR prior for the random effects.
+For each simulated dataset, we fit two Bayesian Poisson GLMMs, the first one using an homoscedastic HGP while the second one employing a DAGAR prior for the random effects.
 Both specified the fixed effects correctly with the single covariate.
 The priors for the parameters in the HGP model were set to be those presented in Section 3.2.
-The priors for the parameters in the DAGAR model were set to the same as those in(Datta et al.,, 2019).
-In particular, we set $\pi(\beta_{0})\propto 1$ and $\beta_{1}\sim N(0,1000^{2})$ for the regression parameters; a Gamma prior with shape 2 and rate 1 was placed on the conditional precision ($\phi$), and a uniform prior between 0 and 1 on the spatial dependence parameter ($\psi$).
-The GoF was evaluated using the criteria in Section 3.2, namely, LOOIC and WAIC, to compare which model fits the data better.
+The priors for the parameters in the DAGAR model were set to the same as those in (datta2019spatial); For more details, see the Supplementary Material, Section S.4.1.
+The GoF was evaluated using the criteria in Section 3.2.
+In the Supplementary Material, Section S.4.2, we also evaluated the HGP’s ability to recover true model parameters under correct specification.
 Figure 2: ECDFs displaying the GoF comparison between HGP and DAGAR models across different data generating processes (AGP: solid line, BYM: dashed line) and model comparison criteria (columns of the panel) based on 200 simulated datasets.
@@
 The figure displays empirical cumulative distribution functions (ECDFs) of the differences ($\Delta_{G}$) in LOOIC and WAIC (left and right columns of the panel, respectively) between the models.
-While no formal thresholds exist for significant differences in the model comparison criteria we used, we consider differences exceeding 5 to be noteworthy, following a similar practice used with the DIC (Vranckx et al.,, 2021).
+While no formal thresholds exist for significant differences in the model comparison criteria we used, we consider differences exceeding 5 to be noteworthy, following a similar practice used with the DIC (vranckx2021stability).
 For AGP-generated data (solid line), our simulation study provides strong evidence that the HGP outperforms the DAGAR in GoF.
@@
 In contrast, evidence favoring DAGAR ($\Delta_{\textrm{LOOIC}}>4$) was observed in only 1% of the generated data.
-Interestingly, the HGP remained competitive even when we generated data from the BYM model (a structure closer to the DAGAR).
+Interestingly, the HGP remained competitive when we generated data from the BYM model (a structure closer to the DAGAR).
 While there was some evidence favoring the DAGAR, $\Delta_{\textrm{LOOIC}}$ was below 4 for more than 50% of the simulations under this scenario.
 These results highlight the HGP’s robustness as a disease mapping model, demonstrating superior performance when the data aligns with its structure and remains a viable option even when a specialized competitor better represents the underlying process.
-In the Appendix Material (Table S.1), we also evaluated the HGP’s ability to recover true model parameters under correct specification.
-We assessed estimation accuracy using Mean Absolute Percentage Error (MAPE) and the frequentist coverage percentage (CP) of 95% credible intervals.
-All model parameters achieved CPs close to the nominal level, with the lowest being 92.5% for $\rho$, which consistently proved the most challenging to estimate.
-The marginal standard deviation ($\sigma$) exhibited a maximum MAPE of only 7%.
 4.2 Data Fusion
-Figure 4: Maps utilized in the data fusion simulation study.
-The left map displays the region used for fitting the models, while the right map shows the data points used to evaluate the models’ predictive accuracy.
 To assess the predictive performance of the HGP against standard data fusion models, we conducted a simulation study mirroring usual data fusion scenarios.
 Within a square region with an area of $51^{2}$, we generated a map comprising a $17\times 17$ grid-boxes of areal units (each with an area of $3^{2}$) and 35 point-referenced data points.
-Additionally, we created five extra spatial units for each of the following three types exclusively to evaluate predictions: larger polygons (grid boxes with area of $3^{2}$), smaller polygons (grid boxes with area of $1.5^{2}$), and point-referenced, as displayed in Figure 4.
-We carefully chose the study region and grid box dimensions so each polygon’s area exceeded one, aligning with typical data fusion problems.
+Additionally, we created five extra spatial units for each of the following three types exclusively to evaluate predictions: larger polygons (grid boxes with area of $3^{2}$), smaller polygons (grid boxes with area of $1.5^{2}$), and point-referenced, as displayed in Figure S.6 from the Supplementary Material.
+We carefully chose the study region and grid box dimensions so each polygon’s area exceeded one.
 This design emulates scenarios where areal units, often derived from satellite or computer models, are more abundant than point-referenced data.
+Moreover, the grid-boxes of different sizes used to assess the predictions have the purpose of investigating how the predictions deteriorate by decreasing the size of the spatial units.
 Hausdorff distances between sample units in this map ranged from $0.4$ to $69.8$.
@@
 This section’s spatial GLMMs employ two distinct spatial random effect structures: a heteroscedastic HGP and an AGP.
-The HGP’s SD function $v(\mathbf{s})$ is defined as $\exp\{\alpha_{0}+\alpha_{1}\mathbbm{1}(\lvert\mathbf{s}\rvert>0)\}$, allowing for distinct variances between point-referenced and areal data.
-We selected HGP parameters ($\tau^{2}=0.17^{2}$, $\sigma=4.4416$, $\sigma_{a}=1.3114$, $\rho=28.45$, and $\nu=0.8$) to closely resemble the data analysis presented in Section 5.2.
+The HGP’s SD function $v(\mathbf{s})$ is defined as $\exp\{\alpha_{0}+\alpha_{1}\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)\}$, allowing for distinct variances between point-referenced and areal data.
+While the AGP is a zero-mean second-order stationary GP with an isotropic Matérn covariance function with smoothness parameter $\kappa=1$.
+See the Appendix for further details on the AGP model, and moraga2017geostatistical for a detailed definition of the Matérn covariance function.
+We selected HGP parameters ($\tau^{2}=0.17^{2}$, $\sigma=4.4416$, $\sigma_{a}=1.3114$, $\rho=28.45$, and $\nu=0.8$) to mirror what was estimated in the data analysis presented in Section 5.2.
 Note that the spatial variance for point and areal units ($\sigma$ and $\sigma_{a}$, respectively) is a one-to-one transformation of $(\alpha_{0},\alpha_{1})$$^{\top}$.
@@
 This value differs from our estimate in the real data due to a necessary rescaling based on the map used in this simulation study.
-For further details, refer to Section S.3 in the Appendix Material.
-The AGP model assumes:
-[…]
-where $\{Z(\mathbf{s})\,:\,\mathbf{s}\in D\}$ is a zero-mean, second-order stationary GP with an isotropic Matérn covariance function with smoothness parameter $\nu=1$.
-SeeMoraga et al., (2017) for a detailed definition of this function.
+For further details, refer to the Supplementary Material, Section S.5.1.
 We simulated AGP data with parameters $\sigma^{2}=2.71^{2}$, $\rho=2.29$, and $\tau^{2}=1\times{10}^{-16}$, again mirroring the real data analysis in Section 5.2.
 In aggregated models, $\tau$ approaches zero as only the average of homoscedastic independent residuals is observed, which converges in probability to zero.
-We then generated 200 datasets from the Normal GLMM under each of the two different spatial random effects’ structures detailed next.
-For each simulated dataset, we fit two Bayesian Normal GLMMs, one using the aforementioned heteroscedastic HGP prior for the random effect and the other using the AGP prior.
+We then generated 200 datasets from the Normal GLMM under each of the two different spatial random effects’ structures detailed above.
+For each simulated dataset, we fit three Bayesian Normal GLMMs for each simulated dataset, one using the aforementioned heteroscedastic HGP prior for the random effect and other two using the AGP prior.
 When fitting the HGP model, we used the priors highlighted in Section 3.2.
-In addition, we put a penalized complexity prior (Simpson et al.,, 2017) on $\tau^{-2}$, such that $\mathbbm{P}(\tau>a)=.05$, where $a$ is half the empirical SD observed in the data.
+In addition, we put a penalized complexity prior (simpson2017penalising) on $\tau^{-2}$, such that $\mathbbm{P}(\tau>a)=.05$, where $a$ is half the empirical SD observed in the data.
 Inferences and predictions were based in 2000 MCMC samples with a 1000 warm-up phase per chain with four parallel chains.
-The initialization of the algorithm is described in Section 3.2.
-Inferences for the model induced by the spatial random effect from Equation (9) were performed using the integrated nested Laplace approximation and the stochastic partial differential equation (Lindgren et al.,, 2011, SPDE) method.
-The joint distribution induced by this AGP involve approximating stochastic integrals.
-Practitioners must define a grid with a fixed number of points within each areal unit to approximate these integrals (Wang and Furrer,, 2021).
-The SPDE uses a mesh to discretize the space to achieve such an approximation efficiently.
-The number of grid boxes per polygon will depend on the resolution of the mesh.
-We used two different meshes to show the inferences and, consequently, the predictions may change substantially depending on the accuracy of the approximation for the stochastic integrals in Equation (9).
-In the Appendix Material, Figure S.2 visualizes these meshes: a sparse mesh (left column), and a fine-resolution mesh (second column), with approximately 2 and 25 points per polygon, respectively.
-Henceforth, we will differentiate the AGP models by mesh density, denoting the sparse-mesh model as AGP1 and the fine-mesh model as AGP2.
-Their priors are detailed in Section S.4 of the Appendix Material.
-We assess the quality of model predictions across different scenarios using four metrics.
+Inferences for GLMMs with an AGP model are detailed in Section S.5.3 from the Supplementary Material.
+An important aspect of the AGP model lies in the approximation of the stochastic integrals’ from Equation (6).
+They are usually approximated numerically (fuentes2005model).
+In order to show the inferences and the predictions may change substantially depending on the accuracy of these approximations, we considered a two different grids to compute those integrals numerically: a sparse- (denoted, AGP1) and a fine-resolution (denoted, AGP2) grid, with approximately 2 and 25 points per polygon, respectively.
+Their priors are detailed in Section S.5.3 of the Supplementary Material.
+In Supplementary Material, Section S.5.4, we also present the results for parameter estimation under the correct specification of the the HGP model.
+We assess the quality of model predictions across different scenarios using three metrics.
 Firstly, point and interval predictions are derived from the posterior predictive distributions (Equation (5)), using their median and the 2.5% and 97.5% percentiles, respectively.
 The root mean square error of prediction (RMSP) is used to evaluate point estimates.
-The remaining three metrics assess interval predictions: the frequentist coverage percentage of prediction intervals, denoted CPP, their width, and the interval score (Gneiting and Raftery,, 2007, IS).
+The remaining two metrics assess interval predictions: the frequentist coverage percentage of prediction intervals, denoted CPP, and the interval score (gneiting2007strictly, IS).
 The latter is defined as
@@
 Table 1: Out-of-sample prediction for the three models under different data generating models.
-“Sp.
-unit” indicates the geometry of the sample units for which predictions were assessed; for polygons (denoted Poly), area is additionally specified within parenthesis.
 Fit specifies the model fitted to the data.
-RMSP is the root mean squared error of prediction, CPP is the frequentist coverage percentage of the 95% prediction interval, Width is the prediction interval width, and IS is the interval score.
+RMSP is the root mean squared error of prediction, CPP is the frequentist coverage percentage of the 95% prediction interval, and IS is the interval score.
 Double headers denote the true data-generating model.
 AGP HGP
-Sp. unit Fit RMSP CPP Width IS RMSP CPP Width IS
-Overall AGP1 1.56 67.8 3.10 17.74 1.81 57.9 2.09 31.37
-AGP2 1.39 95.7 5.59 6.60 1.80 88.1 5.53 10.63
-HGP 1.65 90.8 6.10 8.65 1.51 95.7 6.15 7.05
-Poly ($3^{2}$) AGP1 0.78 75.8 1.91 5.79 0.89 72.5 1.94 6.96
-AGP2 0.60 96.3 2.50 2.87 0.80 89.7 2.68 4.14
-HGP 0.83 97.9 3.78 4.00 0.74 96.4 3.00 3.36
-Poly ($1.5^{2}$) AGP1 1.33 75.6 3.56 11.73 0.90 75.6 2.12 6.47
-AGP2 1.13 95.8 4.69 5.38 1.09 96.6 4.72 5.32
-HGP 1.50 79.2 3.84 9.36 0.85 94.8 3.36 4.01
-Points AGP1 2.56 52.2 3.85 35.70 3.64 22.8 2.20 80.70
-AGP2 2.43 95.1 9.59 11.54 3.51 77.9 9.19 22.42
-HGP 2.62 95.3 10.7 12.57 2.94 96.0 12.09 13.79
-Table 1 presents the prediction assessment metrics under both data-generating models at different types of spatial units.
-Let us first consider the “overall” case, that is, ignoring the types.
+Fit RMSP CPP IS RMSP CPP IS
+AGP1 1.56 67.8 17.74 1.81 57.9 31.37
+AGP2 1.39 95.7 6.60 1.80 88.1 10.63
+HGP 1.65 90.8 8.65 1.51 95.7 7.05
+Table 1 presents the prediction assessment metrics under both data-generating models.
 When the AGP is the true model, the fine mesh AGP model yields the most accurate point predictions (lowest RMSP).
 However, the HGP model’s performance is only marginally worse, with a RMSP 18% higher.
-For interval predictions, the HGP model produces prediction intervals (PI) with near-nominal CPP and widths comparable to the AGP2 model, resulting in a competitive interval score (IS).
-Despite having the narrowest PIs on average, the sparse-mesh AGP model suffers from poor coverage and, consequently, a high IS.
+For interval predictions, the HGP model produces prediction intervals (PI) with near-nominal CPP and competitive IS.
+The AGP1 model suffers from poor coverage and, consequently, a high IS.
 As expected, when the true data-generating model is the HGP, the HGP model’s predictive performance surpasses that of its competitors.
@@
 In contrast, our model eliminates the need for arbitrary mesh selection and consistently delivers highly competitive predictions, enhancing prediction accuracy.
-Table 1 further reveals a consistent decline in prediction accuracy for all models as the area of the spatial units decreased.
-For AGP-generated data, the AGP2 model’s RMSP for point-referenced data was approximately four times higher than for the polygons with bigger area (i.e., $3^{2}$).
-Similarly, the HGP model’s RMSP increased by a factor of 3.15 for the same comparison (point-referenced vs largest polygon).
-For both the HGP and AGP2, CPP remained near the nominal level for all the spatial units.
-However, HGP’s CPP dropped below nominal level for smaller polygons (area of $1.5^{2}$).
-This is likely due to variance underestimation: while the AGP-generated data’s variance is inversely proportional to spatial units, the fitted HGP model only distinguish between points and polygons, not their area.
-Regarding the IS, both our model and the AGP2 model showed IS values roughly three times higher for point-referenced data compared to the largest polygons.
-This aligns with the simulation parameters.
-The AGP1 model’s IS performance deteriorated substantially when decreasing the area of the spatial units.
-Under the HGP data-generating process, the HGP model consistently outperformed competitors on all assessment metrics except PI width.
-Although point predictions were broadly similar across models, the HGP provided remarkably superior interval predictions for all spatial unit types, mantaining CPP near the nominal level (see Table 1).
-In contrast, the AGP1 model exhibited a CPP of only 22.8% for point-referenced data.
-The poor coverage for both AGP models with point-referenced data is further reflected in their IS, which were 5.9 (AGP1) and 1.6 (AGP2) times higher than the HGP.
-On the other hand, the AGP2 model remained competitive with the HGP for polygons interval predictions, while the sparse AGP performed poorly.
-Table S.3 in the Appendix Material presents the results for parameter estimation under the correct specification of the the HGP model.
-We used the same metrics as in Section 4.1 to assess the parameters estimation.
-The MAPE for $\rho$ was around $10.9\%$ while for $\sigma_{a}$ it was $9.21\%$.
-Similarly to what is observed with geostatistical models, the small scale standard deviation parameter, namely $\tau$, was the most challenging parameter to be estimated.
-Unlike the areal simulation, the coverage of some of the variance parameters was consistently below the nominal level of $95\%$.
-This results are expected, as the HGP inherits the weak identifiability of these parameters from geostatistical models under a normal likelihood (Zhang,, 2004).
-Predictions remain trustworthy despite this issue.
+Additional analyses assessing predictions across different spatial resolutions are presented in the Supplementary Material, Section S.5.4.
 5 Applications
@@
 We first analyze a dataset on the count of hospital admissions due to respiratory disease in 2010 for the 134 intermediate zones (IZ) to the north of the river Clyde in the Great Glasgow and Clyde health board in Scotland.
-This dataset is publicly available in theCARBayesdata package (Lee,, 2022).
+This dataset is publicly available in the CARBayesdata package (lee2022carbayesdata).
 For each IZ, the expected number of admissions based on the national age- and sex-standardized rates, alongside the percentage of people classified as income deprived in each IZ are also available.
-The standardized morbidity ratio (SMR), the ratio between the number of observed and expected cases, is often used to control for confounding factors when estimating risks (Greenland,, 1984).
+The standardized morbidity ratio (SMR), the ratio between the number of observed and expected cases, is often used to control for confounding factors when estimating risks (greenland1984bias).
 The goals of the analyses are: (1) to provide smooth estimates of the SMR, adjusted for the percentage of income deprivation, and (2) to contrast the outcomes from the proposed methodology with other established methods.
@@
 We analyze the data using the GLMM model from Section 3.1, with a Poisson likelihood, a $\log$ link function, an intercept, a covariate $x_{i}$, and an offset $E_{i}$.
-We employed the homoscedastic HGP (with a PEC function with smoothness $\nu=0.7$) as the prior for the random effects; the choice of $\nu=0.7$ is justified in Table S.4 of the Appendix Material.
-The priors for the parameters in the HGP GLMM were set to be those presented in Section 3.2.
+We employed the homoscedastic HGP (with a PEC function with smoothness $\nu=0.7$) as the prior for the random effects; the choice of $\nu=0.7$ is justified in Table S.4 of the Supplementary Material.
+The priors and initialization scheme for the parameters in the HGP GLMM match those presented in Section 3.2.
 Inference was performed using four parallel chains with 5000 MCMC samples, thinned by 10, after discarding the initial 1000 iterations as the warm-up period.
-The initialization of the algorithm is described in Section 3.2.
 The highest split-$\hat{R}$ was smaller than $1.01$, providing evidence of convergence.
-Two competing models, BYM and DAGAR, were considered as the priors for the spatial random effects $\mathbf{z}$ for comparison, as outlined in Sections 4.1 and S.2.1, respectively.
-In the BYM model, parametrized by the marginal precision (Riebler et al.,, 2016), we followed Morris et al., (2019) and set standard normal priors for the intercept ($\beta_{0}$), regression coefficient ($\beta_{1}$), and the $\mathrm{logit}$ of the mixing parameter $\xi$.
+Two competing models, BYM and DAGAR (see the Appendix for further details), were considered as the priors for the spatial random effects $\mathbf{z}$ for comparison.
+In the BYM model, parametrized by the marginal precision (riebler2016intuitive), we followed morris2019Bayesian and set standard normal priors for the intercept ($\beta_{0}$), regression coefficient ($\beta_{1}$), and the $\mathrm{logit}$ of the mixing parameter $\xi$.
 Similarly, a standard normal distribution truncated on positive real numbers was used for $\sigma$.
-The priors for the DAGAR model are the ones presented in Section 4.1.
+The priors for the DAGAR model are the ones presented in Section S.4.1.
 The initialization of the MCMC algorithm followed the same principle used for the HGP model.
@@
 For these models, the parameters split-$\hat{R}$ provided evidence of convergence.
-In addition, trace and density plots for posterior samples of all fitted model parameters are provided in Section S.5.1 of the Appendix Material.
+In addition, trace and density plots for posterior samples of all fitted model parameters are provided in Section S.6.1 of the Supplementary Material.
 Table 2: Parameter estimates and model comparison criteria for the three models fitted to the areal dataset. Marginal posterior summary through median and 95% credible intervals.
@@
 When considering the coefficient linked to the covariate representing income deprivation, all three models yield similar posterior estimates.
-The positive posterior estimates the respective parameter ($\beta_{1}$) indicate a positive association between proportion of income deprived people and hospitalization rates for respiratory diseases.
-Among these models, DAGAR exhibits the broadest credible intervals for the intercept, whereas the marginal posterior distribution for this parameter in the BYM model is highly concentrated around 0.22-0.22.
-Examining the spatial aspect, the confidence interval for the practical range parameter $\rho$ in the HGP model implies that the distance at which the spatial correlation between random effects from two distinct spatial units decreases to $0.10$ is at most $6.95$ kilometers.
-Conversely, the 95% CI for the spatial dependence parameter $\psi$ from the DAGAR model spans from $0.07$ to $0.83$.
-In other words, this credible interval suggests that the clustering effect ranges from almost negligible to very strong.
-A similar pattern is observed for the parameter $\xi$ in the BYM model.
-Figure 5: Density histogram of the pairwise Hausdorff distances observed in the dataset and. In blue, estimated posterior density of the practical range $\rho$, and, in red, the prior used for $\rho$.
-To gain a better understanding of this value in the context of the data, Figure 5 displays a density histogram of all pairwise Hausdorff distances observed in the dataset.
+All models found a positive association between income deprivation and hospitalizations ($\beta_{1}$).
+However, the HGP provided a more precise and interpretable estimate of spatial dependence via its practical range parameter, whereas the credible intervals for the dependence parameters in the BYM and DAGAR models were very wide, suggesting inconclusive evidence of spatial clustering.
+To gain a better understanding of this value in the context of the data, in the Supplementary Material, Figure S.9 displays a density histogram of all pairwise Hausdorff distances observed in the dataset.
 The plot also features the prior distribution (in red) and estimated posterior density (in blue) associated with the practical range.
 Notably, the majority of the posterior distribution mass for $\rho$ falls below the concentration of most pairwise distances, indicating a negligible clustering effect in hospital admissions.
-An additional analysis of the spatial correlation is presented in Section S.5 in the Appendix Material.
+An additional analysis of the spatial correlation is presented in Section S.6 in the Supplementary Material.
 In Table 2, we also present two model comparison criteria described in Section 4 for the models fitted to the data.
@@
 However, none of the models delivered a remarkably better fit to this dataset.
-This is further reinforced by looking at the maps of observed and smoothed SMR estimates presented in Figure S.3 from the Appendix Material.
+This is further reinforced by looking at the maps of observed and smoothed SMR estimates presented in Figure S.8 from the Supplementary Material.
 5.2 Data fusion: Air pollution in Ventura and Los Angeles
-Figure 6: PM2.5 in Ventura and Los Angeles counties in California, United States. Points represent direct measurements. Tiles are satellite derived estimates.
 Our second application analyzes fused fine particulate matter (PM2.5) data in Ventura and Los Angeles counties, California, USA.
-The dataset combines point-referenced direct measurements (U.S.
-Environmental Protection Agency Clean Air Markets Division,, 2018) from 1999 to present with areal satellite-derived estimates (Van Donkelaar et al.,, 2015), averaged from 2010 to 2012 (Figure 6).
+The dataset combines point-referenced direct measurements (epa2022clean) with areal satellite-derived estimates (van2015global), averaged from 2010 to 2012 (Supplementary Material, Figure S.23).
 To align temporally, we averaged the direct measurements over the same period.
@@
 Building on previous concepts, we introduce the notation for this application.
-The response variable is the recorded PM2.5, noted $Y(\mathbf{s}_{i})$ at the spatial unit $\mathbf{s}_{i}$, which is either a point or a grid box.
-We again resort to the GLMM model from Section 3.1.
-In this case, there are no covariates, the likelihood is Normal with variance $\tau^{2}$, and the link function is the identity.
+The response variable is the recorded PM2.5, denoted $Y(\mathbf{s}_{i})$.
+In this case, $\mathbf{s}_{i}$ is either a point or a grid box.
+We again resort to the GLMM model from Section 3.1, but without any covariates.
+We assume the likelihood is Normal with variance $\tau^{2}$, and the link function is the identity.
 A heteroscedastic HGP prior is placed on the random effect allowing for differing variances between point-referenced and areal data through the SD function from Equation (2) with $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$.
 For interpretability, we present results for $\sigma=\exp\{\alpha_{0}\}$ and $\sigma_{a}=\exp\{\alpha_{0}+\alpha_{1}\}$ instead of $\alpha_{0}$ and $\alpha_{1}$.
-The model was fit with $\nu$ fixed at different values, and we selected the model with best goodness-of-fit ($\nu=0.8$).
-For more details, see Table S.5 in the Appendix Material.
+The model was fit with $\nu$ fixed at different values, and we selected the one with best goodness-of-fit ($\nu=0.7$).
+For more details, see Table S.5 in the Supplementary Material.
 The priors for model parameters are as in Section 4.2.
-The MCMC algorithm initialization is described in Section 3.2.
-We used four parallel chains, each comprising a warm-up period of 2000 samples and an additional 2000 samples for inference.
-All model parameters exhibited split-$\hat{R}$ values below, indicating good convergence.
-This is further supported by the trace plots presented in Section S.6.1 of the Appendix Material.
-For comparison, we also fit an AGP random effect model, with priors for the model parameters as described in Section 4.2.
-Two meshes of different resolutions were considered; see Figure S.17 in the Appendix Material.
-The sparse- and fine-mesh AGP models are denoted AGP1 and AGP2, respectively.
+We used four parallel chains, each comprising a warm-up period of 1000 samples and an additional 1000 samples for inference.
+All model parameters exhibited split-$\hat{R}$ values below 1, indicating good convergence.
+This is further supported by the trace plots presented in Section S.7.1 of the Supplementary Material.
+For comparison, we also fit two models with AGP random effects, with priors for the model parameters as described in Section 4.2.
+The two models only differ in the precision of the numerical integrals approximation.
+The sparse- and fine-resolution AGP models are denoted AGP1 and AGP2, respectively.
 Table 3: Parameter estimates and out-of-sample prediction assessment for the three models fitted to the fused dataset.
 Marginal posterior summary through median and 95% credible intervals.
-RMSP is the root mean squared error of prediction; CPP is the frequentist coverage of the 95% prediction interval; Width is the width of the prediction interval; and IS is the interval score.
+RMSP is the root mean squared error of prediction; CPP is the frequentist coverage of the 95% prediction interval; and IS is the interval score.
 HGP $\rm AGP_{1}$} $\rm AGP_{2}$}
-$\beta$ 5.605 (4.688, 6.449) 6.219 (2.16, 10.10) 6.187 (5.879, 6.478)
-$\rho$ 13.826 (7.817, 23.608) 13.163 (5.140, 30.244) 0.632 (0.456, 0.832)
-$\tau$ 0.178 (0.073, 0.312) 1.394 (1.242, 1.563) 0.539 (0.394, 0.719)
-$\sigma$ 3.849 (2.921, 5.181) 1.703 (0.960, 2.910) 2.399 (2.024, 2.837)
-$\sigma_{a}$ 1.241 (1.037, 1.505)
-RMSP 1.05 1.45 1.64
-Width 3.57 2.53 9.67
-CPP 95.5 78.6 95.5
-IS 4.80 15.11 13.65
+$\beta$ 5.593 (4.659, 6.431) 6.176 (2.160, 9.990) 6.344 (6.102, 6.585)
+$\rho$ 15.076 (8.445, 28.561) 13.093 (4.809, 29.720) 0.493 (0.340, 0.750)
+$\tau$ 0.174 (0.078; 0.324) 1.403 (1.253; 1.573) 0.539 (0.394, 0.719)
+$\sigma$ 3.893 (2.962; 5.359) 1.717 (0.933; 2.938) 2.877 (2.144; 3.566)
+$\sigma_{a}$ 1.275 (1.064; 1.590)
+RMSP 2.22 3.55 4.45
+CPP 95.6 77.3 93.6
+IS 5.08 15.4 9.55
 We present the posterior median and 95% CI of the model parameters in Table 3.
 When looking at the common parameters between the three models (namely, $\beta_{0}$, $\sigma$, $\rho$, and $\tau$), the HGP acts as a compromise between AGP1 and AGP2 while eliminating the arbitrary choice of the level of discretization for approximating numerical integrals.
-Unlike the aggregated models, our method allows the data to tell us whether the variance of the underlying spatial parameter is smaller or larger than the same quantity for point-referenced data.
-Interestingly, the estimated SD of the point-referenced data is 3.09 times larger then the areal data (95% CI: 2.43, 4.12).
-The interpretation of $\rho$, presented in tens of kilometers, in both models is carried out in a similar fashion because the Hausdorff distance between points is equivalent to the distance between points used by the AGP models.
-Table 3 also presents the out-of-sample prediction performance of the three models, assessed via 10-fold cross-validation.
-Using four metrics from our prior simulation study (RMSP, CPP, prediction interval width, IS), we find the HGP offers the most reliable point predictions, evidenced by its lowest RMSP.
+Unlike the aggregated models, our method allows the data to tell us whether the variance associated to areal units is smaller or larger than that of point-referenced data.
+Interestingly, the estimated SD of the point-referenced data is 3.09 times larger than the areal data (95% CI: 2.43, 4.12).
+The interpretation of $\rho$, presented in tens of kilometers, in both models is carried out in a similar fashion because the Hausdorff distance reduces to the Euclidean distance between points at the point-level.
+Additionally, Table 3 presents the out-of-sample prediction performance of the three models, assessed via 10-fold cross-validation.
+Using the three metrics from our prior simulation study (RMSP, CPP, IS), we find the HGP offers the most reliable point predictions, evidenced by its lowest RMSP.
 The HGP also maintains nominal coverage for its prediction intervals.
-While the AGP2 achieves similar coverage, its prediction intervals are considerably wider, resulting in inferior interval scores.
-Table S.6 in the Appendix Material provides a breakdown by data type, highlighting the poor performance of competing methods for point-referenced data.
+While the AGP2 achieves similar coverage, its interval scores are considerably higher.
+Table S.6 in the Supplementary Material provides a breakdown by data type, highlighting the poor performance of competing methods for point-referenced data.
 The PI frequentist coverage, for example, can be as low as 3%.
-The drastic differences in the AGP results for different mesh resolutions are striking.
+In Table S.7 of the Supplementary Material, we further compared the predictive performance of those models to two additional alternatives: a Log-Normal model with HGP random effects, and a GP where the areal units are reduced to their centroids, resulting in a geostatistical model.
+We have not tested an AGP where the response follows a Log-Normal distribution, because the random effects would need to be aggregated on the response scale (suen2025coherent), which can be seen as a consequence of Jensen’s inequality.
+Surprisingly, the AGP models performed the worst, while the Log-Normal HGP provided the second-best predictive performance.
+The results from the AGP models diverged significantly depending on the mesh resolution.
 The credible interval width for the intercept varies significantly between the two models.
-Crucially, the practical range reveals a key divergence: AGP1 indicates moderate spatial dependence, while AGP2 suggests weak dependence, likely explaining the observed overfitting in the latter.
+The AGP1 model indicates moderate spatial dependence, while AGP2 suggests weak dependence.
 This weak dependence may imply AGP2 assigns nearly independent random intercepts to each data point, with variances inversely proportional to the sample unit’s area.
-AGP mesh resolution is typically practitioner-informed, and results are highly sensitive to this tuning parameter, which lacks a consensus on “optimal resolution” (Fuentes and Raftery,, 2005; Liu et al.,, 2011; Wang and Furrer,, 2021).
+AGP mesh resolution is typically practitioner-informed, and results are highly sensitive to this tuning parameter, which lacks a consensus on “optimal resolution” (fuentes2005model; liu2011empirical; wang2021combining).
 Figure 3: Change-of-support of PM2.5 to a finer resolution using the sparse and fine AGP and the HGP priors.
@@
 The proposed framework unifies the modeling of spatial data, extending geostatistical models to include areal and fused data within the same setting.
-As a prior for spatial random effects in a GLMM, the HGP proves to be competitive to specialized models for diverse spatial data types.
+Despite the mathematical simplicity, the subtle change of paradigm powered by the HGP opens a new direction of research in spatial statistics while taking advantage of what has been proposed in the geostatistics and GP literature.
+The proposal has the potential to simplify how we deal with complex problems such as spatial misalignment and change of support.
+In simpler scenarios, as a prior for spatial random effects in a GLMM, the HGP proves to be competitive to models specially designed for areal and fused data types.
 Our simulation studies indicate that the HGP is particularly advantageous for areal data with units of varying sizes and shapes, as confirmed by our disease mapping application.
-Unlike its specialized counterparts, the HGP also provides a directly interpretable spatial parameter, offering insights into the practical range of spatial dependence.
+Unlike the traditional areal models, the HGP also provides a directly interpretable spatial parameter, offering insights into the practical range of spatial dependence.
 In data fusion contexts, the HGP delivers more reliable interval predictions compared to the sparse AGP, and it provided similar results to the fine AGP in parameter estimation and spatial range.
@@
 This success is even more remarkable considering the HGP bypasses the subjective step of defining a grid or mesh, a crucial but problematic requirement for the AGP.
-In addition, independent of the spatial data type, it continues to have direct and interpretable parameters.
-Our proposed methodology shares the “big nn” computational complexity challenge inherent in traditional geostatistical approaches.
-Consequently, the computational cost of evaluating the finite-dimensional likelihood induced by the HGP for a set of $n$ areal units scales as $\mathcal{O}(n^{3})$.
-Proving the validity of the PEC function under the Hausdorff distance on Euclidean space was a demanding theoretical task.
-Relaxing the isotropy assumption for this function presents further theoretical complexities.
-Moreover, to apply the HGP to data measured on manifolds (like the sphere), new theoretical results are needed to guarantee the positive definiteness of a correlation function within that specific geometric context.
-The HGP framework offers numerous opportunities for development.
-Further researching its theoretical properties will strengthen the foundation for its applications.
-A crucial focus should be developing correlation functions designed specifically for the HGP, particularly those suited to scenarios where spatial data is observed on a sphere.
-Approximations inspired by methods like the Nearest Neighbor Gaussian Process (Datta et al.,, 2016; Quiroz et al.,, 2023) or reduced rank approximations (Solin and Särkkä,, 2020) could significantly broaden the applicability of the HGP to large-scale datasets.
-These advancements hold the potential to establish the HGP as an even more powerful and versatile tool for spatial data analysis.
+The empirical assessment in Section 2.3 reveals an interesting insight: the properties of the induced process appear linked to the geometric complexity of the sets involved.
+This is most evident in the contrasting results from our two scenarios.
+For the data fusion scenario, which involves only convex sets, all correlation matrices were positive definite.
+In contrast, for the areal data scenario, where most spatial units are non-convex, we observed negative eigenvalues at the border of the parameter space (Figure 1).
+This contrast provides strong empirical evidence that the PEC function is a valid and robust choice for applications involving convex sets, such as most data fusion problems.
+While our extensive empirical assessments consistently demonstrate the positive definiteness of the PEC function across practical parameter ranges, a formal proof establishing the general conditions under which functions of the Hausdorff distance are positive definite remains an open and challenging theoretical question.
+Nevertheless, the proposed HGP addresses several long-standing problems in spatial statistics.
+For instance, traditional areal models are unsuitable for out-of-sample prediction and often have counter-intuitive parameters interpretation (wall2004close; assuncao2009neighborhood).
+Furthermore, the validity of CAR models, for instance, also depends on constraining the parametric space to ensure a non-singular precision matrix (wall2004close).
+While data fusion models do not share the same issues, we have shown that their results can be significantly affected by the arbitrary precision of required approximations.
+By introducing the HGP and demonstrating its utility, we aim to provide a foundation for future theoretical work, thereby bridging the gap between practical application and complete mathematical understanding.
+Several conceptual challenges for the HGP remain open for future investigation.
+Relaxing the isotropy assumption, for instance, requires further theoretical development.
+To apply the HGP to data on manifolds like the sphere, new results are needed to guarantee the process’s validity.
+Furthermore, because the HGP is a set-valued function, any analysis of its curvature (e.g., halder2024Bayesian) necessitates more general concepts of differentiation (khastan2021new).
+It is important to note, however, that this latter complexity does not prevent the use of powerful inference frameworks that rely on the posterior’s gradient, such as variational inference (blei2017variational).
+Although the $\mathcal{O}(n^{3})$ computational complexity is a limitation, the method was developed with the intent of leveraging what has already been developed for GPs.
+Thus, adapting scalable nearest neighbor approximations (datta2016hierarchical; quiroz2023fast) for the HGP is a viable avenue for future research.
+The main challenge would be to identify nearest neighbors in the Hausdorff distance sense while avoiding the storage of a $n\times n$ distance matrix.
+Conflict of interest
+All authors declare that they have no conflicts of interest.
+Appendix A Review of Existing Methods
+The BYM model we use corresponds to a zero-mean multivariate Normal model with covariance matrix defined as
+$\sigma^{-2}((1-\xi)\bm{I}+\xi\bm{Q}^{-}),$
+where $\sigma$ is the average marginal SD, $\bm{I}$ is the identity matrix, and $\bm{Q}^{-}$ is the generalized inverse of the scaled precision matrix of an ICAR model.
+The spatial dependence parameter $\xi$ is a mixing parameter.
+Its parametric space comprises of real numbers between 0 and 1, with larger values indicate stronger spatial dependence.
+The entries of $\bm{Q}$ are as follows:
+$Q_{ij}=\begin{cases}kn_{N(i)},&\text{ if }i=j,\\ -k,&\text{ if }i\sim j,\\ 0,&\text{ elsewhere,}\end{cases}$
+where $n_{N(i)}$ denotes the number of neighbors for the region $i$, and $i\sim j$ indicates that $j$ and $i$ are neighbors, and $k$ is the scaling constant.
+The scaling constant depends on the adjacency matrix and enhances the interpretation of $\sigma$, which becomes the average marginal standard deviation, as opposed to the conditional standard deviation in the original formulation (riebler2016intuitive).
+We used the BYM as a prior for the random effects in the spatial GLMMs (Section 3.1) for areal data in Section 5.1.
+We also generated data from a spatial GLMM assuming the BYM was the actual distribution of the spatial random effects for part of the simulation studies in Section 4.1.
+The DAGAR model, in our context, is a zero-mean multivariate Normal model.
+We used the DAGAR as a prior for the spatial GLMM (Section 3.1) random effects (e.g., $\mathbf{z}$) for areal data, in particular, in Section 4.1 and 5.1.
+The precision matrix for the DAGAR model is constructed using the data ordering, an adjacency matrix, and two parameters: a conditional precision, $\tau$, and a spatial dependence parameter, $\psi$.
+The $\psi$ parameter represents the average spatial correlation between adjacent spatial units.
+Specifically, the data ordering and adjacency matrix define a tree graph, and the precision matrix is derived from an order-1 autoregressive process on the nodes of this graph.
+A key advantage of this model is its computational efficiency, as it can be evaluated using either closed-form conditional distributions or direct formulas for the Cholesky decomposition of its covariance matrix.
+In particular, the Cholesky decomposition of the precision matrix under a DAGAR model can be written as $\mathbf{F}^{-1/2}(\mathbf{I}-\mathbf{B})$, where $\mathbf{F}$ is a diagonal matrix, $\mathbf{I}$ is the identify matrix, and $\mathbf{B}$ is lower-triangular (datta2022nearest).
+The entries of $\mathbf{F}$ and $\mathbf{B}$, respectively, are given by
+$F_{ii}$ $={[\tau(1+1+n_{\vec{N}(i)}\psi^{2})]}^{-1}\quad\text{ and }$
+$B_{ij}$ $=\mathbbm{1}(j\sim i,j<i)\frac{\psi}{1+1+n_{\vec{N}(i)}\psi^{2}},$
+where $\mathbbm{1}(A)$ denotes an indicator function, equaling $1$ if the event $A$ happens and 0 otherwise.
+Note that, in this case, $n_{\vec{N}(i)}=\sum_{j}\mathbbm{1}(j\sim i,j<i)$ represents the number of “directional neighbors”.
+Although the DAGAR depends on an arbitrary order for the data points, the authors also developed an “order-free” variant of the model.
+However, the “order-free” DAGAR does not enjoy the computational efficiency of its ordered counterpart.
+The AGP model is employed as a prior on the random effects ($\mathbf{z}$) of the spatial GLMM (Section 3.1) for fused data.
+In particular, we used this model to generate data for the simulation studies in Section 4, and to analyze the dataset from Section 5.2.
+The AGP core assumption is as follows:
+[…]
+where $\{Z(\mathbf{s})\,:\,\mathbf{s}\in D\}$ is a (point-level) zero-mean, second order stationary and isotropic GP (cressie1993statistics, p.
+53) defined over the study region, $\mathcal{A}(\cdot)$ is a function that returns the area of a spatial object.